### v1

# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

Abstract
Designing efficient algorithms for multi-agent reinforcement learning (MARL) is fundamentally challenging due to the fact that the size of the joint state and action spaces are exponentially large in the number of agents. These difficulties are exacerbated when balancing sequential global decision-making with local agent interactions. In this work, we propose a new algorithm SUBSAMPLE-MFQ (Subsample-Mean-Field-Q-learning) and a decentralized randomized policy for a system with $n$ agents. For $k\leq n$, our algorithm system learns a policy for the system in time polynomial in $k$. We show that this learned policy converges to the optimal policy in the order of $\widetilde{O}(1/\sqrt{k})$ as the number of subsampled agents $k$ increases. We validate our method empirically on Gaussian squeeze and global exploration settings.
1 Introduction
Reinforcement Learning (RL) has become a popular learning framework to solve sequential decision making problems in unknown environments, and has achieved tremendous success in a wide array of domains such as playing the game of Go (Silver et al., 2016), robotic control (Kober et al., 2013), and autonomous driving (Kiran et al., 2022; Lin et al., 2023). A critical feature of most real-world systems is their uncertain nature, and consequently RL has emerged as a powerful tool for learning optimal policies for multi-agent systems to operate in unknown environments (Kim & Giannakis, 2017; Zhang et al., 2021; Lin et al., 2024; Anand & Qu, 2024). While the early literature on RL predominantly focused on the single-agent setting, multi-agent reinforcement learning (MARL) has also recently achieved impressive successes in a broad range of areas, such as coordination of robotic swarms (Preiss et al., 2017), self-driving vehicles (DeWeese & Qu, 2024), real-time bidding (Jin et al., 2018), ride-sharing (Li et al., 2019), and stochastic games (Jin et al., 2020).
Despite growing interest in multi-agent RL (MARL), extending RL to multi-agent settings poses significant computational challenges due to the curse of dimensionality (Sayin et al., 2021). Even if the individual agents’ state or action spaces are small, the global state space or action space can take values from a set with size that is exponentially large as a function of the number of agents. For example, even model-free RL algorithms such as temporal difference (TD) learning (Sutton et al., 1999) or tabular $Q$-learning require computing and storing a $Q$-function (Bertsekas & Tsitsiklis, 1996) that is as large as the state-action space. Unfortunately, in MARL, the joint state-action space is exponentially large in the number of agents. In the case where the system’s rewards are not discounted, reinforcement learning on multi-agent systems is provably NP-hard (Blondel & Tsitsiklis, 2000), and such scalability issues have been observed in the literature in a variety of settings (Guestrin et al., 2003; Papadimitriou & Tsitsiklis, 1999; Littman, 1994). Independent $Q$-learning (Tan, 1997) seeks to overcome these scalability challenges by considering other agents as a part of the environment; however, this often fails to capture a key feature of MARL, which is inter-agent interactions.
Even in the fully cooperative regime, MARL is fundamentally difficult, since agents in the real-world not only interact with the environment but also with each other (Shapley, 1953). An exciting line of work that addresses this intractability is mean-field MARL (Lasry & Lions, 2007; Yang et al., 2018; Gu et al., 2021, 2022b, 2022a; Hu et al., 2023). The mean-field approach assumes that all the agents are homogeneous in their state and action spaces, enabling their interactions to be approximated by a two-agent setting: here, each agent interacts with a representative “mean agent” which is the empirical distribution of states of all other agents. Under these homogeneity assumptions, mean-field MARL allows learning of optimal policies with a sample complexity that is polynomial in the number of agents. However, when the number of homogeneous agents is large, storing a polynomially-large $Q$ table (where the polynomial’s degree depends on the size of the state space for a single agent) can still be infeasible. This motivates the following fundamental question:
Can we design an efficient and approximately optimal MARL algorithm for policy-learning in a cooperative multi-agent system with many agents?
Contributions.
We answer this question affirmatively. Our key contributions are outlined below.
  • Subsampling Algorithm. We propose SUBSAMPLE-MFQ, an algorithm designed to address the challenge of MARL with a large number of local agents. We model the problem as a Markov Decision Process (MDP) with a global agent and $n$ local agents. SUBSAMPLE-MFQ selects $k\leq n$ local agents to learn a deterministic policy $\widehat{\pi}^{\mathrm{est}}_{k}$, by applying mean-field value iteration on a subsystem with $k$ local agents to learn $\widehat{Q}_{k}^{\mathrm{est}}$, which can be viewed as a smaller $Q$ function. It then deploys a stochastic policy $\widehat{\pi}_{k}$ which works as follows: the global agent uniformly samples $k$ local agents at each step and uses $\widehat{\pi}_{k}$ to determine its action, whereas each local agent uniformly samples $k-1$ other local agents and uses $\widehat{\pi}_{k}$ to determine its action.
  • Sample Complexity and Theoretical Guarantee. As the number of local agents increases, the size of $\widehat{Q}_{k}$ scales polynomially with $k$, rather than polynomially with $n$ as in mean-field MARL. Analogously, when the size of the local agent’s state space grows, the size of $\widehat{Q}_{k}$ scales exponentially with $k$, rather than exponentially with $n$, as in traditional $Q$-learning). The key analysis technique underlying our results is a novel MDP sampling result. Through it, we show that the performance gap between ${{\pi}_{k}^{\mathrm{est}}}$ and the optimal policy ${\pi^{*}}$ is $\widetilde{O}(1/\sqrt{k})$. The choice of $k$ reveals a fundamental trade-off between the size of the $Q$-table and the optimality of ${\pi}_{k}^{\mathrm{est}}$. When $k\!=\!O(\log n)$, SUBSAMPLE-MFQ is the first MARL algorithm to achieve a run-time that is polylogarithmic in $n$, representing an exponential speedup over the previously best-known polytime mean-field MARL methods, while maintaining a decaying optimality gap.
  • Numerical Simulations. We evaluate the effectiveness of SUBSAMPLE-MFQ in two scenarios: a bounding box problem, and a Gaussian squeeze problem. Our experiments reveal a monotonic improvement in the learned policies as $k\to n$, providing a substantial speedup over mean-field $Q$-learning.
While our results are theoretical in nature, it is our hope that SUBSAMPLE-MFQ will lead to further exploration into the potential of subsampling in general stochastic/Markovian games and networked multi-agent reinforcement learning, and inspire the development of practical algorithms for multi-agent settings.
1.1 Related Literature
MARL has a rich history, starting with early works on Markov games used to characterize the decision-making process (Littman, 1994; Sutton et al., 1999), which can be regarded as a multi-agent extension of the Markov Decision Process (MDP). MARL has since been actively studied (Zhang et al., 2021) in a broad range of settings. MARL is most similar to the category of “succinctly described” MDPs (Blondel & Tsitsiklis, 2000), where the state/action space is a product space formed by the individual state/action spaces of multiple agents, and where the agents interact to maximize an objective. A promising line of research that has emerged over recent years constrains the problem to sparse networked instances to enforce local interactions between agents (Qu et al., 2020a; Lin et al., 2020; Mondal et al., 2022). In this formulation, the agents correspond to vertices on a graph who only interact with nearby agents. By exploiting Gamarnik’s correlation decay property from combinatorial optimization (Gamarnik et al., 2009), they overcome the curse of dimensionality by simplifying the problem to only search over the policy space derived from the truncated graph to learn approximately optimal solutions. However, as the underlying network structure becomes dense with many local interactions, the neighborhood of each agent gets large, causing these algorithms to become intractable.
Mean-Field RL. Under assumptions of homogeneity in the state/action spaces of the agents, the problem of densely networked multi-agent RL was answered in Yang et al. (2018); Gu et al. (2021) by approximating the solution with a mean-field approach where the approximation error scales in $O(1/\sqrt{n})$. To avoid designing algorithms on probability spaces, they study MARL under Pareto optimality and use the law of large numbers to consider a lifted space with a mean agent that aggregates the system’s rewards and dynamics. It then applies kernel regression on $\epsilon$-nets of the lifted space to design policies in time polynomial in $n$. In contrast, our work achieves subpolynomial runtimes by directly sampling from this mean-field distribution. Cui & Koeppl (2022) introduce heterogeneity to mean-field MARL by modeling non-uniform interactions through graphons; however, these methods make critical assumptions on the existence of a sequence of graphons converging in cut-norm to the finite instance. In the cooperative setting, Cui et al. (2023) considers a mean-field setting with $q$ types of homogeneous agents; however, their approach does not converge to the optimum policy.
Structured RL. Our work is related to factored MDPs and exogenous MDPs. In factored MDPs, there is a global action affecting every agent whereas in our case, each agent has its own action (Min et al., 2023; Lauer & Riedmiller, 2000). Our result has a similar flavor to MDPs with exogenous inputs from learning theory, (Dietterich et al., 2018; Foster et al., 2022; Anand & Qu, 2024), where our subsampling algorithm treats each sampled state as an endogenous state, but where the exogenous dependencies can be dynamic.
Miscellaneous. Our work adds to the growing literature on the Centralized Training with Decentralized Execution paradigm (Wang et al., 2023; Zhou et al., 2023), as our algorithm learns a provably approximately optimal policy using centralized information, but makes decisions using local information during execution. In the distributed setting, recent works on V-learning (Jin et al., 2020) reduce the dependence of the product action space to an additive dependence. In contrast, our work further accomplishes a reduction on the complexity of the joint state space, which has not been previously accomplished for dense graphs. Finally, one can approximate the large $Q$-table through linear function approximation (Jin et al., 2020) which significantly reduces the computational complexity. However, achieving theoretical bounds on the performance loss caused by the function approximation is generally intractable without making stronger assumptions such as Linear Bellman completeness (Golowich & Moitra, 2024) or low Bellman-Eluder dimension (Jin et al., 2021a). While our work primarily consider a tabular setting with finite spaces, we provide a further extension of it to the non-tabular setting, under assumptions of Linear Bellman completeness.
2 Preliminaries
In this section, we formally introduce the problem, state some examples for our setting, and provide technical details of the mean-field and $Q$-learning techniques that will be used throughout the paper.
Notation.
For $k,n\in\mathbb{N}$ where $k\leq n$, let $\binom{[n]}{k}$ denote the set of $k$-sized subsets of $[n]=\{1,\dots,n\}$. For any vector $z\in\mathbb{R}^{d}$, let $\|z\|_{1}$ and $\|z\|_{\infty}$ denote the standard $\ell_{1}$ and $\ell_{\infty}$ norms of $z$ respectively. Let $\|\mathbf{A}\|_{1}$ denote the matrix $\ell_{1}$-norm of $\mathbf{A}\in\mathbb{R}^{n\times m}$. Given a collection of variables $s_{1},\dots,s_{n}$ the shorthand $s_{\Delta}$ denotes the set $\{s_{i}:i\in\Delta\}$ for $\Delta\subseteq[n]$. We use $\widetilde{O}(\cdot)$ to suppress polylogarithmic factors in all problem parameters except $n$. For a discrete measurable space $(\mathcal{X},\mathcal{F})$, the total variation distance between probability measures $\rho_{1},\rho_{2}$ is given by $\mathrm{TV}(\rho_{1},\rho_{2})=\frac{1}{2}\sum_{x\in\mathcal{X}}|\rho_{1}(x)-\rho_{2}(x)|$. Next, $x\sim\mathcal{D}(\cdot)$ denotes that $x$ is a random element sampled from a probability distribution $\mathcal{D}$, and we denote the uniform distribution over a finite set $\Omega$ by $\mathcal{U}(\Omega)$.
2.1 Problem Formulation
We consider a system of $n+1$ agents, where agent $g$ is a “global decision making agent” and the remaining $n$ agents, denoted by $[n]$, are “local agents.” At time step $t$, the agents are in state $s(t)=(s_{g}(t),s_{1}(t),...,s_{n}(t))\in\mathcal{S}\coloneq\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, where $s_{g}(t)\in\mathcal{S}_{g}$ denotes the global agent’s state, and for each $i\in[n]$, $s_{i}(t)\in\mathcal{S}_{l}$ denotes the state of the $i$’th local agent. The agents cooperatively select actions $a(t)=(a_{g}(t),a_{1}(t),...,a_{n}(t))\in\mathcal{A}$ where $a_{g}(t)\in\mathcal{A}_{g}$ denotes the global agent’s action and $a_{i}(t)\in\mathcal{A}_{l}$ denotes the $i$’th local agent’s action. At each time-step $t$, the next state for all the agents is independently generated by stochastic transition kernels $P_{g}:\mathcal{S}_{g}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\to[0,1]$ and $P_{l}:\mathcal{S}_{l}\times\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\to[0,1]$ as follows:
$s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),$ (1)
$s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t)),\forall i\in[n].$ (2)
The system then collects a structured stage reward $r(s(t),a(t))$ where the reward $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ depends on $s(t)$ and $a(t)$ through Eq. 3, and where the choice of functions $r_{g}$ and $r_{l}$ is typically application specific.
$r(s,a)=\underbrace{r_{g}(s_{g},a_{g})}_{{\text{global component}}}+\frac{1}{n}\sum_{i\in[n]}\underbrace{r_{l}(s_{i},s_{g},a_{i})}_{\text{local component}},\qquad\forall(s,a)\in\mathcal{S}\times\mathcal{A}.$ (3)
We define a policy $\pi$ as a mapping from $\mathcal{S}$ to $\mathcal{A}$ where we want the policy $\pi$ to maximize the value function for each $s\in S$, which is the expected discounted reward $V^{\pi}(s)=\mathbb{E}_{a(t)\sim\pi(\cdot|s)}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s\right]$, where $\gamma\in(0,1)$ is a discounting factor.
Notably, the cardinality of the search space simplex for the optimal policy is $|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n}$, which is exponential in the number of agents. When noting that the local agents are all homogeneous, and therefore permutation-invariant with respect to the rewards of the system (the order of the other agents does not matter to any single decision-making agent), techniques from mean-field MARL restrict the cardinality of the search space simplex for the optimal policy to $|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|n^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}$, reducing the exponential complexity on $n$ to a polynomial complexity on $n$. In practical systems, when $n$ is large, the $\mathrm{poly}(n)$ sample-complexity may still be computationally infeasible. Therefore, the goal of this problem is to learn an approximately optimal policy with subpolynomial sample complexity, further overcoming the curse of dimensionality.
Example 2.1 (Gaussian Squeeze).
In this task, $n$ homogeneous agents determine their individual action $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, $a_{i}=\{0,\dots,9\}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. In scenarios of traffic congestion, each agent $i\in[n]$ is a traffic controller trying to send $a_{i}$ vehicles into the main road, where controllers coordinate with each other to avoid congestion, hence avoiding either over-use or under-use, thereby contributing to the entire system. This GS problem serves as an ablation study on the impact of subsampling for MARL.
Example 2.2 (Constrained Exploration).
Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent wishes to constrain the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect stage rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, by only using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where some shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Therefore, through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while avoiding unnecessary risk or inefficiency.
Figure 1: Constrained exploration for warehouse accidents.
Figure 2: Traffic congestion settings with Gaussian squeeze.
Due to space constraints, we leave the details of the experiments to Appendix J.
To efficiently learn policies that maximize the objective, we make the following standard assumptions:
Assumption 2.3 (Finite state/action spaces).
We assume that the state and action spaces of all the agents in the MARL game are finite: $|\mathcal{S}_{l}|,|\mathcal{S}_{g}|,|\mathcal{A}_{g}|,|\mathcal{A}_{l}|<\infty$. Appendix H weakens this assumption to the non-tabular setting of infinite compact sets.
Assumption 2.4 (Bounded rewards).
The global and local components of the reward function are bounded. Specifically, $\|r_{g}(\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{g}$, and $\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{l}$. This implies that $\|r(\cdot,\cdot)\|_{\infty}\leq\widetilde{r}_{g}+\widetilde{r}_{l}:=\widetilde{r}$.
Definition 2.5 ($\epsilon$-optimal policy).
Given an objective function $V$ and policy simplex $\Pi$, a policy $\pi\in\Pi$ is $\epsilon$-optimal if $V(\pi)\geq\sup_{\pi^{*}\in\Pi}V(\pi^{*})-\epsilon$.
Remark 2.6.
Heterogeneity among the local agents can be captured by modeling agent types as part of the agent state: assign a type $\varepsilon_{i}\in\mathcal{E}$ to each local agent $i\in[n]$ by letting $\mathcal{S}_{l}=\mathcal{E}\times\mathcal{S}^{\prime}_{l}$, where $\mathcal{E}$ is a set of possible types that are treated as a fixed part of the agent’s state. The transition and reward functions can vary depending on the agent’s type. The global agent can provide unique signals to local agents of each type by letting $s_{g}\in\mathcal{S}_{g}$ and $a_{g}\in\mathcal{A}_{g}$ denote a state/action vector where each element matches to a type $\varepsilon\in\mathcal{E}$.
2.2 Technical Background
Q-learning.
To provide background for the analysis in this paper, we review a few key technical concepts in RL. At the core of the standard Q-learning framework (Watkins & Dayan, 1992) for offline-RL is the $Q$-function $Q:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. $Q$-learning seeks to produce a policy $\pi^{*}(\cdot|s)$ that maximizes the expected infinite horizon discounted reward. For any policy $\pi$, $Q^{\pi}(s,a)=\mathbb{E}^{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s,a(0)=a]$. One approach to learning the optimal policy $\pi^{*}(\cdot|s)$ is dynamic programming, where the $Q$-function is iteratively updated using value-iteration: $Q^{0}(s,a)=0$, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Then, for all $t\in[T]$, $Q^{t+1}(s,a)=\mathcal{T}Q^{t}(s,a)$, where $\mathcal{T}$ is the Bellman operator defined as
$\mathcal{T}Q^{t}(s,a)=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}}Q^{t}(s^{\prime},a^{\prime}).$ (4)
The Bellman operator $\mathcal{T}$ is $\gamma$-contractive, which ensures the existence of a unique fixed-point $Q^{*}$ such that $\mathcal{T}Q^{*}=Q^{*}$, by the Banach fixed-point theorem (Banach, 1922). Here, the optimal policy is the deterministic greedy policy $\pi^{*}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}\to\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$, where $\pi^{*}(s)=\arg\max_{a\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}}Q^{*}(s,a)$. However, the complexity of a single update to the $Q$-function is $O(|\mathcal{S}_{g}||\mathcal{S}_{l}|^{n}|\mathcal{A}_{g}||\mathcal{A}_{l}|^{n})$, which grows exponentially with $n$. As the number of local agents increases $(n\boldsymbol{\mathit{g}}|\mathcal{S}_{l}|$), this exponential update complexity renders $Q$-learning impractical.
Mean-field Transformation.
To address this, mean-field MARL (under homogeneity assumptions) studies the distribution function $F_{z_{[n]}}:\mathcal{Z}_{l}\to\mathbb{R}$, where $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$, given by
$F_{z_{[n]}}(z)\coloneq\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{s_{i}=z_{s},a_{i}=z_{a}\},\quad\forall z\coloneq(z_{s},z_{a})\in\mathcal{S}_{l}\times\mathcal{A}_{l}.$ (5)
Let $\mu_{n}(\mathcal{Z}_{l})=\{\frac{b}{n}|b\in\{0,\dots,n\}\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$ be the space of $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-sized tables, where each entry is an element of $\{0,\frac{1}{n},\frac{2}{n},\dots,1\}$. In this space, $F_{z_{[n]}}\in\mu_{n}{(\mathcal{Z}_{l})}$ where $F_{z_{[n]}}$ represents the proportion of agents in each state/action pair. The $Q$-function is permutation-invariant in the local agents, since permuting the labels of homogeneous local agents with the same state will not change the action of the decision-making agent. Hence, $Q(s_{g},s_{[n]},a_{g},a_{[n]})=\widehat{Q}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})$. Here, $\widehat{Q}:\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mathcal{A}_{g}\times\mathcal{A}_{1}\times\mu_{n-1}(\mathcal{Z}_{l})\to\mathbb{R}$ is a reparameterized $Q$-function learned by mean-field value iteration: one initializes $\widehat{Q}^{0}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=0$. At each time-step $t$, we update $\widehat{Q}$ as $\widehat{Q}^{t+1}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=\widehat{\mathcal{T}}\widehat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})$, where $\widehat{\mathcal{T}}$ is the Bellman operator in distribution space:
$\widehat{\mathcal{T}}\widehat{Q}^{t}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in[n]\end{subarray}}\max_{\begin{subarray}{c}(a_{g}^{\prime},a_{1}^{\prime},a_{[n]\setminus 1}^{\prime})\\ \in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mathcal{A}_{l}^{n-1}\end{subarray}}\widehat{Q}^{t}(s^{\prime}_{g},s^{\prime}_{1},a^{\prime}_{g},a^{\prime}_{1},F_{z^{\prime}_{[n]\setminus 1}}).$ (6)
Since $\mathcal{T}$ is a $\gamma$-contraction, so is $\widehat{\mathcal{T}}$. Hence, $\widehat{T}$ has a unique fixed-point $\widehat{Q}^{*}$ such that $\widehat{Q}^{*}(s_{g},s_{1},a_{g},a_{1},F_{z_{[n]\setminus 1}})=Q^{*}(s_{g},s_{[n]},a_{g},a_{[n]})$ and the optimal policy is the deterministic greedy policy given by $\widehat{\pi}^{*}(s_{g},s_{1},F_{s_{[n]\setminus 1}})=\arg\max_{(a_{g},a_{1},a_{[n]\setminus 1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mathcal{A}_{l}^{n-1}}\widehat{Q}^{*}(s_{g},s_{1},a_{g},a_{1},F_{s_{[n]\setminus 1},a_{[n]\setminus 1}},a_{g})$. The update complexity to the $\widehat{Q}$-function is $O(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|})$, which scales polynomially in $n$.
Remark 2.7.
The solution offered by mean-field value iteration and standard $Q$-learning requires a sample complexity of $\min\{\widetilde{O}(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|^{n}),\widetilde{O}(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|})\}$, where one uses standard $Q$-learning when $|\mathcal{Z}_{l}|^{n-1}<n^{|\mathcal{Z}_{l}|}$, and mean-field value iteration otherwise. In each of these regimes, as $n$ scales, the update complexity can become incredibly computationally intensive. Therefore, we introduce the SUBSAMPLE-MFQ algorithm to mitigate the cost of scaling the number of local agents.
3 Method and Theoretical Results
In this section, we propose the SUBSAMPLE-MFQ algorithm to overcome the polynomial (in $n$) sample complexity of mean-field value iteration and the exponential (in $n$) sample complexity of traditional $Q$-learning. In our algorithm, the global agent randomly samples a subset of local agents $\Delta\subseteq[n]$ such that $|\Delta|=k$, for $k\leq n$. It ignores all other local agents $[n]\setminus\Delta$, and performs value iteration to learn the $Q$-function $\widehat{Q}_{k,m}^{*}$ and policy $\widehat{\pi}_{k,m}^{*}$ for this surrogate subsystem of $k$ local agents, where $m$ is the number of samples used to update the $Q$-functions’ estimates of the unknown system. Here, When $|\mathcal{Z}_{l}|^{k-1}<k^{|\mathcal{Z}_{l}|}$, the algorithm uses traditional value-iteration (Algorithm 1), and when $|\mathcal{Z}_{l}|^{k-1}>k^{|\mathcal{Z}_{l}|}$, it uses mean-field value iteration (Algorithm 2). The surrogate reward gained by this subsystem at each time step is $r_{\Delta}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$:
$r_{\Delta}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{g},s_{i},a_{i}).$ (7)
To convert the optimality of each agent’s action within the $k$ local-agent subsystem to an approximate optimality on the full $n$-agent system, we use a randomized policy ${\pi}^{\mathrm{est}}_{k,m}$ (Algorithm 3), where the global agent samples $\Delta\in\mathcal{U}\binom{[n]}{k}$ at each time-step to derive the action $a_{g}\leftarrow\widehat{\pi}_{k,m}^{*}(s_{g},s_{\Delta})$, and where each local $i$ agent samples $k-1$ other local agents $\Delta_{i}$ to derive the action $\widehat{\pi}_{k,m}^{*}(s_{g},s_{i},s_{\Delta_{i}})$. Finally, Theorem 3.4 shows that the policy ${\pi}_{k,m}^{\mathrm{est}}$ converges to the optimal policy $\pi^{*}$ as $k\to n$. We first present Algorithms 2 and 1 (SUBSAMPLE-MFQ: Learning) and Algorithm 3 (SUBSAMPLE-MFQ: Execution), which we describe below. For this, a crucial characterization is the notion of the empirical distribution function:
Definition 3.1 (Empirical Distribution Function).
For any population $(z_{1},\dots,z_{n})\in\mathcal{Z}_{l}^{n}$, where $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$, define the empirical distribution function $F_{z_{\Delta}}:\mathcal{Z}_{l}\to\mathbb{R}_{+}$ for $\Delta\subseteq[n]$ such that $|\Delta|=k$ by:
$F_{z_{\Delta}}(x)\coloneq F_{s_{\Delta},a_{\Delta}}(x)\coloneq\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{s_{i}=z_{s},a_{i}=z_{a}\},\quad\forall z\coloneq(z_{s},z_{a})\in\mathcal{S}_{l}\times\mathcal{A}_{l}.$ (8)
Let $\mu_{k}(\mathcal{Z}_{l})\!\coloneq\!\left\{\frac{b}{k}|b\in\{0,\dots,k\}\right\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$ be the space of $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-length vectors where each entry in a vector is an element of $\{0,\frac{1}{k},\frac{2}{k},\dots,1\}$ such that $F_{z_{\Delta}}\in\mu_{k}(\mathcal{Z}_{l})$. Here, $F_{z_{\Delta}}$ is the proportion of agents in the $k$-local-agent subsystem at each state.
Algorithms 2 and 1 (Offline learning). Let $m\in\mathbb{N}$ denote the sample size for the learning algorithm with sampling parameter $k\leq n$. When $|\mathcal{Z}_{l}|^{k-1}\!\leq\!k^{|\mathcal{Z}_{l}|}$, we empirically learn the optimal $Q$-function for a subsystem with $k$-local agents denoted by $\widehat{Q}_{k,m}^{\mathrm{est}}\!:\!\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}\to\mathbb{R}$: set $\widehat{Q}_{k,m}^{0}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!0$ for all $(s_{g},s_{\Delta},a_{g},a_{\Delta})\!\in\!\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$. At time step $t$, set $\widehat{Q}_{k,m}^{t+1}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})={\widetilde{\mathcal{T}}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, where ${\widetilde{\mathcal{T}}}_{k,m}$ is the empirically adapted Bellman operator in Eq. 9.
When $|\mathcal{Z}_{l}|^{k-1}\!>\!k^{|\mathcal{Z}_{l}|}$, we empirically learn the optimal mean-field $Q$-function for a $k$ local agent system, denoted (with abuse of notation) by $\widehat{Q}_{k,m}^{\mathrm{est}}\!:\!\mathcal{S}_{g}\!\times\!\mathcal{S}_{1}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}\to\mathbb{R}$. For all $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$, set $\widehat{Q}_{k,m}^{0}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\!=\!0$. At time $t$, set $\widehat{Q}_{k,m}^{t+1}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\!=\!\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$, where $\widehat{\mathcal{T}}_{k,m}$ is the empirically adapted mean-field Bellman operator in Eq. 10.
Since the system is unknown, $\mathcal{T}_{k,m}$ and $\widehat{\mathcal{T}}_{k,m}$ cannot compute the direct expectation from the Bellman operator and instead draw $m$ random samples $s_{g}^{j}\!\sim\!P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\!\sim\!P_{l}(\cdot|s_{i},s_{g},a_{i})$ for each $j\!\in\![m]$, $i\!\in\!\Delta$:
$\begin{split}&{\widetilde{\mathcal{T}}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k,m}^{t}(s_{g}^{j},{s_{\Delta}^{j}},a_{g}^{\prime},a_{\Delta}^{\prime}).\end{split}$ (9)
$\begin{split}&\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{1}^{\prime}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}^{\prime}}\in\mu_{k-1}(\mathcal{A}_{l})}\widehat{Q}_{k,m}^{t}(s_{g}^{j},s_{1}^{j},F_{s_{\widetilde{\Delta}}^{j},a_{\widetilde{\Delta}^{\prime}}},a_{1}^{\prime},a_{g}^{\prime}).\end{split}$ (10)
$\widehat{Q}_{k,m}^{t}$ depends on $s_{\Delta}$ and $a_{\Delta}$ through $F_{z_{\Delta}}$, and $\widetilde{\mathcal{T}}_{k,m}$/$\widehat{\mathcal{T}}_{k,m}$ are $\gamma$-contractive (Lemma B.8). So, Algorithms 1 and 2 apply value iteration with their Bellman operator until $\widehat{Q}_{k,m}$ converges to a fixed point satisfying $\widetilde{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}\!=\!\widehat{Q}_{k,m}^{\mathrm{est}}$ and $\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}\!=\!\widehat{Q}_{k,m}^{\mathrm{est}}$, yielding equivalent deterministic policies $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ and $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\widetilde{\Delta}}})$:
$\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})$ (11)
$\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\widetilde{\Delta}}})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{1}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}}\in\mu_{k-1}(\mathcal{A}_{l})}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$ (12)
Algorithm 3 (Online implementation). Here, Algorithm 3 (SUBSAMPLE-MFQ: Execution) randomly samples $\Delta\!\sim\!\mathcal{U}\binom{[n]}{k}$ at each time step and uses action $a_{g}\!\sim\!\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})$ to get reward $r(s,a_{g})$. This procedure of first sampling $\Delta$ and then applying $\widehat{\pi}^{\mathrm{est}}_{k,m}$ is denoted by a stochastic policy ${\pi}_{k,m}^{\mathrm{est}}(a|s)=[{\pi}_{k,m}^{\mathrm{est}}(a_{g}|s),{\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)]$, where ${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)$ is the global agent’s action distribution and ${\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)$ is the local agent’s action distribution:
${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)=\frac{1}{\binom{n}{k}}\sum_{\Delta\in\binom{[n]}{k}}\mathbbm{1}(\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})=a)$ (13)
${\pi}_{k,m}^{\mathrm{est}}(a_{i}|s)=\frac{1}{\binom{n-1}{k-1}}\sum_{\widetilde{\Delta}\in\binom{[n]\setminus i}{k-1}}\mathbbm{1}(\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})=a_{i})$ (14)
Then, each agent transitions to their next state based on Eq. 1.
Algorithm 1 SUB-SAMPLE-MFQ: Learning (if $|\mathcal{Z}_{l}|^{k-1}\leq k^{|\mathcal{Z}_{l}|}$)
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
1:  Uniformly sample $\Delta\subseteq[n]$ such that $|\Delta|=k$.
2:  Initialize $\widehat{Q}^{0}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0$ for $(s_{g},s_{\Delta},a_{g},a_{\Delta})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$.
3:  for $t=1$ to $T$ do
4:     for $(s_{g},s_{\Delta},a_{g},a_{\Delta})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ do
5:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})=\widetilde{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
6:  Return $\widehat{Q}_{k,m}^{T}$. For all $s_{g},s_{\Delta}\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}$, let $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k,m}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$.
Algorithm 2 SUBSAMPLE-MFQ: Learning (if $|\mathcal{Z}_{l}|^{k-1}<k^{|\mathcal{Z}_{l}|}$)
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
1:  Set $\widetilde{\Delta}=\{2,\dots,k\}$.
2:  Set $\mu_{k-1}(\mathcal{Z}_{l})=\{\frac{b}{k-1}:b\in\{0,1,\dots,k-1\}\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$.
3:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=0$, for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$.
4:  for $t=1$ to $T$ do
5:     for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$ do
6:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$
7:  $\forall(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{S}_{l})$, let $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})=\operatorname*{arg\,max}_{a_{g}\in\mathcal{A}_{g},a_{i}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}}\in\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$.
We define the greedy deterministic policy for the $k$-local-agent subsystem by: $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})$ by:
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta}):=\arg\max_{{a}_{g}^{*},{a}_{\Delta}^{*}}\widehat{Q}_{k,m}^{T}(s_{g},s_{\Delta},{a}_{g}^{*},{a}^{*}_{\Delta}):=([\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g},[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}),$ (15)
where $[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}$ reads the maximizer $a_{g}^{*}$, and $[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}$ reads the maximizer $a_{i}^{*}$. Then, for the $n$-agent system, the global agent samples local agents $\Delta$ uniformly from (${[n]\choose k}$ to derive action $a_{g}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}$, and allows agent $i$ to choose agents $\Delta_{i}$ uniformly from (${[n]\setminus i\choose k-1}$ to derive action $a_{i}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},(s_{\Delta_{i}},s_{i})]_{l}$.
Algorithm 3 SUB-SAMPLE-MFQ: Execution
0:  A multi-agent system as described in Section 2. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},F_{s_{\Delta}})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$, where $s_{0}$ is a distribution on the initial global state $(s_{g},s_{{[n]}})$
2:  Initialize the total reward $R_{0}=0$.
3:  Policy $\pi_{k}^{\mathrm{est}}(s)$ is defined as follows:
4:  for $t=0$ to $T^{\prime}$ do
5:     Choose $\Delta$ uniformly at random from (${[n]\choose k}$ and let $a_{g}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta}(t))]_{g}$.
6:     for $i=1$ to $n$ do
7:        Choose $\Delta_{i}$ uniformly at random from (${[n]\setminus i\choose k-1}$ and let $a_{i}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{i}(t),s_{\Delta_{i}}(t))]_{l}$.
8:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$.
9:     Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
10:     $R_{t+1}=R_{t}+\gamma^{t}\cdot r(s,a)$
3.1 Theoretical Guarantee
This subsection shows that the value of the expected discounted cumulative reward produced by ${\pi}^{\mathrm{est}}_{k,m}$ is approximately optimal, where the optimality gap decays as $k\!\to\!n$ and $m$ becomes large.
Bellman noise. We introduce the notion of Bellman noise, which is used in the main theorem. Consider $\widehat{\mathcal{T}}_{k,m}$. Clearly, it is an unbiased estimator of the generalized adapted Bellman operator $\widehat{\mathcal{T}}_{k}$,
$\begin{split}&\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!r_{\Delta}(s,a)\!+\!\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime}).\end{split}$ (16)
For all $(s_{g},s_{\Delta},a_{g},a_{\Delta})\!\in\!\mathcal{S}_{g}\times\!\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$, set $\widehat{Q}^{0}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0$. For $t\in\mathbb{N}$, let $\widehat{Q}_{k}^{t+1}=\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}$, where $\widehat{\mathcal{T}}_{k}$ is defined for $k\leq n$ in Eq. 16. Then, $\widehat{\mathcal{T}}_{k}$ is also a $\gamma$-contraction (Lemma B.7) with fixed-point $\widehat{Q}_{k}^{*}$. By the law of large numbers, $\lim_{m\to\infty}\widehat{\mathcal{T}}_{k,m}\!=\!\widehat{\mathcal{T}}_{k}$ and $\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\to 0$ as $m\to\infty$. For finite $m$, $\epsilon_{k,m}\coloneq\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}$ is the well-studied Bellman noise.
Lemma 3.2.
By the Chernoff bound, for $k\in[n]$ and $m\in\mathbb{N}$, where $m$ is the number of samples in Eq. 10, $\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\leq\epsilon_{k,m}\leq\widetilde{O}(1/\sqrt{k})$, when $m=m^{*}=\frac{2|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{2.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}}{(1-\gamma)^{5}}\log(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$.
We defer the proof of Lemma 3.2 to Appendix F.1. Next, let ${\pi}_{k}^{\mathrm{est}}\coloneq\widetilde{\pi}_{k,m^{*}}^{\mathrm{est}}$. To compare the difference in the performance of $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$, we define the value function of a policy $\pi$ by $V^{\pi}$:
Definition 3.3.
The value function $V^{\pi}:\mathcal{S}\to\mathbb{R}$ of a given policy $\pi$, for $\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is:
$V^{\pi}(s)=\mathbb{E}_{\begin{subarray}{c}a(t)\sim\pi(\cdot|s(t))\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))\bigg{|}s(0)=s\right].$ (17)
Intuitively, the value function $V^{\pi}(s)$ is the expected discounted cumulative reward when starting from state $s$ and applying actions from the policy $\pi$ across an infinite horizon. With the above preparations, we are primed to present our main result: a decaying bound on the optimality gap for our learned policy ${\pi}_{k}^{\mathrm{est}}$.
Theorem 3.4.
Let $\widetilde{\pi}_{k}$ denote the learned policy from SUBSAMPLE-MFQ. Then, $\forall s\in\mathcal{S}$:
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k}^{\mathrm{est}}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\widetilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{21}{10\sqrt{k}}\leq\widetilde{O}(1/\sqrt{k})$
We provide for the theorem in Appendix C, and defer the proof to Appendix F.
Discussion 3.5.
Between Algorithms 2 and 1, the asymptotic sample complexity to learn $\widehat{\pi}^{\mathrm{est}}_{k}$ for a fixed $k$ is $\min\{O(|\mathcal{Z}_{l}|^{k}),O(k^{|\mathcal{Z}_{l}|})\}$. By Theorem 3.4, as $k\to n$, the optimality gap decays, revealing a fundamental trade-off in the choice of $k$: increasing $k$ improves the performance of the policy, but increases the size of the $Q$-function. We explore this trade-off further in our experiments. For $k=O(\log n)$, the runtime is $\min\{O(n^{\log|\mathcal{Z}_{l}|}),O((\log n)^{|\mathcal{Z}_{l}|})\}$. This is an exponential speedup on the complexity from mean-field value iteration (from $\mathrm{poly}(n)$ to $\mathrm{poly}(\log n)$), as well as over traditional value-iteration (from $\mathrm{exp}(n)$ to $\mathrm{poly}(n)$). Further, the optimality gap decays to 00 at the rate of $O(1/\sqrt{\log n})$.
Appendix G extends the theorem to stochastic reward distributions.
Theorem 3.6.
Suppose we are given two families of distributions on the reward functions, $r_{g}(s_{g},a_{g})\sim\{\mathcal{G}_{s_{g},a_{g}}\}_{s_{g},a_{g}\in\mathcal{S}_{g}\times\mathcal{A}_{g}}$ and $r_{l}(s_{i},s_{g},a_{i})\sim\{\mathcal{L}_{s_{i},s_{g},a_{i}}\}_{s_{i},s_{g},a_{i}\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}$. Then, under standard assumptions of boundedness of the support of $\mathcal{G}_{s_{g},a_{g}}$ and $\mathcal{L}_{s_{i},s_{g},a_{i}}$, SUBSAMPLE-MFQ learns a stochastic policy $\pi_{k}^{\mathrm{est}}$ satisfying
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\pi^{\mathrm{est}}_{k}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}}.$
In the non-tabular setting with infinite state/action spaces, one could replace the $Q$-learning algorithm with any arbitrary value-based RL method that learns $\widehat{Q}_{k}$ with function approximation (Sutton et al., 1999) such as deep $Q$-networks (Silver et al., 2016). Doing so raises an additional error that factors into Theorem 3.4.
Definition 3.7 (Linear MDP).
$\mathrm{MDP}(S,A,\mathbb{P},r)$ is a linear MDP with feature map $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ if there exist $d$ unknown (signed) measures $\mu=(\mu^{1},\dots,\mu^{d})$ over $\mathcal{S}$ and a vector $\theta\in\mathbb{R}^{d}$ such that for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have $\mathbb{P}(\cdot|s,a)=\langle\phi(s,a),\mu(\cdot)\rangle$ and $r(s,a)=\langle\phi(s,a),\theta\rangle$.
Suppose the system is a linear MDP, where $\mathcal{S}_{g}$ and $\mathcal{S}_{l}$ are infinite compact sets. By a reduction from Ren et al. (2024) and using function approximation to learn the spectral features $\phi_{k}$ for $\widehat{Q}_{k}$, we derive a performance guarantee for the learned policy $\pi_{k}^{\mathrm{est}}$, where the optimality gap decays with $k$.
Theorem 3.8.
When ${\pi}^{\mathrm{est}}_{k}$ is derived from the spectral features $\phi_{k}$ learned in $\widehat{Q}_{k}$, and $M$ is the number of samples used in the linear function approximation,
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}^{\mathrm{est}}_{k}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\log\left(2k^{2}\right)\frac{\|{\phi}_{k}\|^{5}}{\sqrt{M}}+\frac{2}{\sqrt{k}}\cdot\frac{\gamma\widetilde{r}}{1-\gamma}\|\bar{\phi}_{k}\|\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\cdot\left(1-\frac{2}{\sqrt{k}}\right)$
We defer the proof of Theorem 3.8 to Appendix H.
Remark 3.9.
If $k=O(\log n)$, SUBSAMPLE-MFQ can handle $|\mathcal{E}|=O(\log n/\log\log n)$ different types of local agents, since the run-time of the learning algorithm becomes $\mathrm{poly}(n)$. This additionally supersedes the previous-best heterogeneity capacity from Mondal et al. (2022), which only handles constant $|\mathcal{E}|$.
Remark 3.10.
Our algorithm and policy also contributes to the growing literature on the centralized-training-decentralized-execution paradigm (Zhou et al., 2023; Wang et al., 2023), and the literature on exogenous MDPs, wherein our algorithm has an advantage in that the agents do not all have to “see” each other during the online (execution) stage, and hence contains a partially observable setting.
4 Conclusion and Future Works
This work develops subsampling for mean field MARL in a cooperative system with a global decision-making agent and $n$ homogeneous local agents. We propose SUBSAMPLE-MFQ which learns each agent’s best response to the mean effect from a sample of its neighbors, allowing an exponential reduction on the sample complexity of approximating a solution to the MDP. We provide a theoretical analysis on the optimality gap of the learned policy, showing that the learned policy converges to the optimal policy with the number of agents $k$ sampled at the rate $\widetilde{O}(1/\sqrt{k})$ validate our theoretical results through numerical experiments. We further extend this result to the non-tabular setting with infinite state and action spaces.
We recognize several future directions. Firstly, this model studies a ‘star-network’ setting to model a single source of density. It would be fascinating to extend this subsampling framework to general networks. We believe expander-graph decompositions (Anand & Umans, 2023; Reingold, 2008) are amenable for this. A second direction would be to find connections between our sub-sampling method to algorithms in federated learning, where the rewards can be stochastic. A third direction of this work would be to consider the setting of truly heterogeneous local agents. Finally, it would be exciting to generalize this work to the online setting without a generative oracle: we conjecture that tools from recent works on stochastic approximation (Chen & Theja Maguluri, 2022) and no-regret RL (Jin et al., 2021b) might be valuable.
Acknowledgements.
We gratefully acknowledge insightful discussions with Siva Theja Maguluri, Sarah Liaw, Yiheng Lin, and Yi Wu.
Appendix A Mathematical Background and Additional Remarks
Outline of the Appendices.
  • Section 5 presents the proof of the Lipschitz continuity between $\widehat{Q}_{k}$ and $Q^{*}$
  • Section 6 presents the TV distance bound
  • Section 7 proves the bound on the optimality gap between the learned policy $\widetilde{\pi}_{k,m}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$
Table 1: Important notations in this paper.
Notation Meaning
$\|\cdot\|_{1}$ $\ell_{1}$ (Manhattan) norm;
$\|\cdot\|_{\infty}$ $\ell_{\infty}$ norm;
$\mathbb{Z}_{+}$ The set of strictly positive integers;
$\mathbb{R}^{d}$ The set of $d$-dimensional reals;
$[m]$ The set $\{1,\dots,m\}$, where $m\in\mathbb{Z}_{+}$;
(${[m]\choose k}$ The set of $k$-sized subsets of $\{1,\dots,m\}$;
$a_{g}$ $a_{g}\in\mathcal{A}_{g}$ is the action of the global agent;
$s_{g}$ $s_{g}\in\mathcal{S}_{g}$ is the state of the global agent;
$a_{1},\dots,a_{n}$ $a_{1},\dots,a_{n}\in\mathcal{A}_{l}^{n}$ are the actions of the local agents $1,\dots,n$
$s_{1},\dots,s_{n}$ $s_{1},\dots,s_{n}\in\mathcal{S}_{l}^{n}$ are the states of the local agents $1,\dots,n$
$a$ $a=(a_{g},a_{1},\dots,a_{n})\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{n}$ is the tuple of actions of all agents
$s$ $s=(s_{g},s_{1},\dots,s_{n})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is the tuple of states of all agents
$z_{i}$ $z_{i}=(s_{i},a_{i})\in\mathcal{Z}_{l}$, for $i\in[n]$
$\mu_{k}(\mathcal{Z}_{l})$ $\mu_{k}(\mathcal{Z}_{l})=\{0,1/k,2/k,\dots,1\}^{|\mathcal{Z}_{l}|}$
$\mu(\mathcal{Z}_{l})$ $\mu(\mathcal{Z}_{l})\coloneq\mu_{n}(\mathcal{Z}_{l})\{0,1/n,2/n,\dots,1\}^{|\mathcal{Z}_{l}|}$
$s_{\Delta}$ For $\Delta\subseteq[n]$, and a collection of variables $\{s_{1},\dots,s_{n}\}$, $s_{\Delta}\coloneq\{s_{i}:i\in\Delta\}$
$\sigma(z_{\Delta},z^{\prime}_{\Delta})$ Product sigma-algebra generated by sequences $z_{\Delta}$ and $z^{\prime}_{\Delta}$;
$\pi^{*}$ $\pi^{*}$ is the optimal deterministic policy function such that $a=\pi^{*}(s)$;
$\widehat{\pi}_{k}^{*}$ $\widehat{\pi}_{k}^{*}$ is the optimal deterministic policy function on a constrained system of $|\Delta|=k$ local agents
$\widetilde{\pi}^{\mathrm{est}}_{k}$ $\widetilde{\pi}^{\mathrm{est}}_{k}$ is the stochastic policy mapping learned with parameter $k$ such that $a\sim\widetilde{\pi}_{k}^{\mathrm{est}}(s)$
$P_{g}(\cdot|s_{g},a_{g})$ $P_{g}(\cdot|s_{g},a_{g})$ is the stochastic transition kernel for the state of the global agent
$P_{l}(\cdot|a_{i},s_{i},s_{g})$ $P_{l}(\cdot|a_{i},s_{i},s_{g})$ is the stochastic transition kernel for the state of any local agent $i\in[n]$
$r_{g}(s_{g},a_{g})$ $r_{g}$ is the global agent’s component of the reward
$r_{l}(s_{i},s_{g},a_{i})$ $r_{l}$ is the component of the reward for local agent $i\in[n]$
$r(s,a)$ $r(s,a)\coloneq r_{[n]}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},s_{g},a_{i})$ is the reward of the system
$r_{\Delta}(s,a)$ $r_{\Delta}(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})$ is the constrained system’s reward with $|\Delta|=k$ local agents
$\mathcal{T}$ $\mathcal{T}$ is the centralized Bellman operator
$\widehat{\mathcal{T}}_{k}$ $\widehat{\mathcal{T}}_{k}$ is the Bellman operator on a constrained system of $|\Delta|=k$ local agents
$\Pi^{\Theta}(y)$ $\ell_{1}$ projection of $y$ onto set $\Theta$;
Definition A.1 (Lipschitz continuity).
Given two metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$ and a constant $L\in\mathbb{R}_{+}$, a mapping $f:\mathcal{X}\to\mathcal{Y}$ is $L$-Lipschitz continuous if for all $x,y\in\mathcal{X}$, $d_{\mathcal{Y}}(f(x),f(y))\leq L\cdot d_{\mathcal{X}}(x,y)$.
Theorem A.2 (Banach-Caccioppoli fixed point theorem (Banach, 1922)).
Consider the metric space $(\mathcal{X},d_{\mathcal{X}})$, and $T:\mathcal{X}\to\mathcal{X}$ such that $T$ is a $\gamma$-Lipschitz continuous mapping for $\gamma\in(0,1)$. Then, by the Banach-Cacciopoli fixed-point theorem, there exists a unique fixed point $x^{*}\in\mathcal{X}$ for which $T(x^{*})=x^{*}$. Additionally, $x^{*}=\lim_{s\to\infty}T^{s}(x_{0})$ for any $x_{0}\in\mathcal{X}$.
Appendix B Notation and Basic Lemmas
For convenience, we restate below the various Bellman operators under consideration.
Definition B.1 (Bellman Operator $\mathcal{T}$).
$\mathcal{T}Q^{t}(s,a):=r(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}Q^{t}(s^{\prime},a^{\prime})$ (18)
Definition B.2 (Adapted Bellman Operator $\widehat{\mathcal{T}}_{k}$).
The adapted Bellman operator updates a smaller $Q$ function (which we denote by $\widehat{Q}_{k}$), for a surrogate system with the global agent and $k\in[n]$ local agents denoted by $\Delta$, using mean-field value iteration and $j\in\Delta$ such that:
$\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\forall i\in\Delta\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}^{t}(s_{g}^{\prime},s_{j}^{\prime},F_{z^{\prime}_{\Delta\setminus j}},a^{\prime}_{j},a^{\prime}_{g})$ (19)
Definition B.3 (Empirical Adapted Bellman Operator $\widehat{\mathcal{T}}_{k,m}$).
The empirical adapted Bellman operator $\widehat{\mathcal{T}}_{k,m}$ empirically estimates the adapted Bellman operator update using mean-field value iteration by drawing $m$ random samples of $s_{g}\sim P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim P_{l}(\cdot|s_{i},s_{g},a_{i})$ for $i\in\Delta$, where for $\ell\in[m]$, the $\ell$’th random sample is given by $s_{g}^{\ell}$ and $s_{\Delta}^{\ell}$, and $j\in\Delta$:
$\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{j},F_{z_{\Delta\setminus j}},a_{j},a_{g}):=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell})$ (20)
Remark B.4.
Since the local agents $1,\dots,n$ are all homogeneous in their state/action spaces, the $\widehat{Q}_{k}$-function only depends on them through their empirical distribution $F_{z_{\Delta}}$. Therefore, throughout the remainder of the paper, we will use $\widehat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})\coloneq\widehat{Q}_{k}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i})\coloneq\widehat{Q}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ interchangeably, unless making a remark about the computational complexity of learning each function.
Lemma B.5.
For any $\Delta\subseteq[n]$ such that $|\Delta|=k$, suppose $0\leq r_{\Delta}(s,a)\leq\widetilde{r}$. Then, for all $t\in\mathbb{N}$, $\widehat{Q}_{k}^{t}\leq\frac{\widetilde{r}}{1-\gamma}$.
Proof.
The proof follows by induction on $t$. The base case follows from $\widehat{Q}_{k}^{0}:=0$. For the induction, note that by the triangle inequality $\|\widehat{Q}_{k}^{t+1}\|_{\infty}\leq\|r_{\Delta}\|_{\infty}+\gamma\|\widehat{Q}_{k}^{t}\|_{\infty}\leq\widetilde{r}+\gamma\frac{\widetilde{r}}{1-\gamma}=\frac{\widetilde{r}}{1-\gamma}$. ∎
Remark B.6.
By the law of large numbers, $\lim_{m\to\infty}\widehat{\mathcal{T}}_{k,m}=\widehat{\mathcal{T}}_{k}$, where the error decays in $O(1/\sqrt{m})$ by the Chernoff bound. Also, $\widehat{\mathcal{T}}_{n}\coloneq\mathcal{T}$. Further, Lemma B.5 is independent of the choice of $k$. Therefore, for $k=n$, this implies an identical bound on $Q^{t}$. An identical argument implies the same bound on $\widehat{Q}_{k,m}^{t}$.
$\mathcal{T}$ satisfies a $\gamma$-contractive property under the infinity norm (Watkins & Dayan, 1992). We similarly show that $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k,m}$ satisfy a $\gamma$-contractive property under infinity norm in Lemmas B.7 and B.8.
Lemma B.7.
$\widehat{\mathcal{T}}_{k}$ satisfies the $\gamma$-contractive property under infinity norm:
$\|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{\prime}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}\|_{\infty}\leq\gamma\|\widehat{Q}_{k}^{\prime}-\widehat{Q}_{k}\|_{\infty}$ (21)
Proof.
Suppose we apply $\widehat{\mathcal{T}}_{k}$ to $\widehat{Q}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ and $\widehat{Q}^{\prime}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ for $|\Delta|=k$. Then:
$\|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{\prime}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}\|_{\infty}$ $=\gamma\max_{\begin{subarray}{c}s_{g}\in\mathcal{S}_{g},\\ a_{g}\in\mathcal{A}_{g},\\ F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})}\end{subarray}}\!\left|\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta,\\ \end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},s_{g},a_{i}),\\ \forall i\in\Delta^{\prime}\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|$
$\leq\gamma\max_{\begin{subarray}{c}s_{g}^{\prime}\in\mathcal{S}_{g},F_{z_{\Delta}}\in\mu_{k}{(\mathcal{Z}_{l})},a^{\prime}\in\mathcal{A}\end{subarray}}\left|\widehat{Q}_{k}^{\prime}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})\right|=\gamma\|\widehat{Q}_{k}^{\prime}-\widehat{Q}_{k}\|_{\infty}$
The equality cancels common $r_{\Delta}(s,a)$ terms in each operator. The second line uses Jensen’s inequality, maximizes over actions, and bounds expected values with the maximizers of the random variables.∎
Lemma B.8.
$\widehat{\mathcal{T}}_{k,m}$ satisfies the $\gamma$-contractive property under infinity norm.
Proof.
Similarly to Lemma B.7, suppose we apply $\widehat{\mathcal{T}}_{k,m}$ to $\widehat{Q}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$ and $\widehat{Q}_{k,m}^{\prime}(s_{g},F_{z_{\Delta}},a_{g})$. Then:
$\|\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k}-\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{\prime}_{k}\|_{\infty}$ $=\frac{\gamma}{m}\left\|\sum_{\ell\in[m]}\left(\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})-\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}^{\prime}_{k}(s_{g}^{\ell},F_{z_{\Delta}^{\ell}},a_{g}^{\prime})\right)\right\|_{\infty}$
$\leq\gamma\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},s_{g}^{\prime}\in\mathcal{S}_{g},z_{\Delta}\in\mathcal{Z}_{l}^{k}\end{subarray}}|\widehat{Q}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})-\widehat{Q}^{\prime}_{k}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})|\leq\gamma\|\widehat{Q}_{k}-\widehat{Q}^{\prime}_{k}\|_{\infty}$
The first inequality uses the triangle inequality and the general property $|\max_{a\in A}f(a)-\max_{b\in A}f(b)|\leq\max_{c\in A}|f(a)-f(b)|$. In the last line, we recover the definition of infinity norm.∎
Remark B.9.
The $\gamma$-contractivity of $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k,m}$ attracts the trajectory between two $\widehat{Q}_{k}$ and $\widehat{Q}_{k,m}$ functions on the same state-action tuple by $\gamma$ at each step. Repeatedly applying the Bellman operators produces a unique fixed-point from the Banach fixed-point theorem which we introduce in Definitions B.10 and B.11.
Definition B.10 ($\widehat{Q}_{k}^{*}$-function).
Suppose $\widehat{Q}_{k}^{0}:=0$ and let $\widehat{Q}_{k}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Denote the fixed-point of $\widehat{\mathcal{T}}_{k}$ by $\widehat{Q}^{*}_{k}$ such that $\widehat{\mathcal{T}}_{k}\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$.
Definition B.11 ($\widehat{Q}_{k,m}^{\mathrm{est}}$-function).
Suppose $\widehat{Q}_{k,m}^{0}:=0$ and let $\widehat{Q}_{k,m}^{t+1}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},F_{z_{\Delta}},a_{g})$ for $t\in\mathbb{N}$. Denote the fixed-point of $\widehat{\mathcal{T}}_{k,m}$ by $\widehat{Q}^{\mathrm{est}}_{k,m}$ such that $\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})=\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})$.
Corollary B.11.1.
Observe that by backpropagating results of the $\gamma$-contractive property for $T$ time steps:
$\|\widehat{{Q}}_{k}^{*}-\widehat{{Q}}_{k}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\widehat{Q}_{k}^{*}-\widehat{Q}_{k}^{0}\|_{\infty},\quad\quad\quad\|\widehat{{Q}}_{k,m}^{\mathrm{est}}-\widehat{{Q}}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\cdot\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}^{0}_{k,m}\|_{\infty}$ (22)
Further, noting that $\widehat{Q}_{k}^{0}=\widehat{Q}_{k,m}^{0}:=0$, $\|\widehat{Q}_{k}^{*}\|_{\infty}\leq\frac{\widetilde{r}}{1-\gamma}$, and $\|\widehat{Q}_{k,m}^{\mathrm{est}}\|_{\infty}\leq\frac{\widetilde{r}}{1-\gamma}$ from Lemma B.5:
$\|\widehat{Q}_{k}^{*}-\widehat{Q}_{k}^{T}\|_{\infty}\leq\gamma^{T}\frac{\widetilde{r}}{1-\gamma},\quad\quad\quad\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k,m}^{T}\|_{\infty}\leq\gamma^{T}\frac{\widetilde{r}}{1-\gamma}$ (23)
Remark B.12.
Corollary B.11.1 characterizes the error decay between $\widehat{Q}_{k}^{T}$ and $\widehat{Q}_{k}^{*}$ and shows that it decays exponentially in the number of Bellman iterations by a $\gamma^{T}$ multiplicative factor.
Furthermore, we characterize the maximal policies greedy policies obtained from $Q^{*},\widehat{Q}_{k}^{*}$, and $\widehat{Q}_{k,m}^{\mathrm{est}}$.
Definition B.13 (Optimal policy $\pi^{*}$).
The greedy policy derived from $Q^{*}$ is
$\pi^{*}(s):=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a).$ (24)
Definition B.14 (Optimal subsampled policy $\widehat{\pi}_{k}^{*}$).
The greedy policy from $\widehat{Q}_{k}^{*}$ is
$\widehat{\pi}_{k}^{*}(s_{g},s_{i},F_{s_{\Delta\setminus i}})\coloneq\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k}^{*}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (25)
Definition B.15 (Optimal empirically subsampled policy $\widehat{\pi}_{k,m}^{\mathrm{est}}$).
The greedy policy from $\widehat{Q}_{k,m}^{\mathrm{est}}$ is given by
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}}):=\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{i},a_{g}).$ (26)
Figure 4 details the analytic flow on how we use the empirical adapted Bellman operator to perform value iteration on $\widehat{Q}_{k,m}$ to get $\widehat{Q}_{k,m}^{\mathrm{est}}$ which approximates $Q^{*}$.
Algorithm 4 gives a stable implementation of Algorithm 2 with learning rates $\{\eta_{t}\}_{t\in[T]}$. Algorithm 4 is provably numerical stable under fixed-point arithmetic (Anand et al., 2024). $\widehat{Q}_{k,m}^{t}$ in Algorithm 4 follows the $\gamma$-contractivity as in Lemma B.7, given an appropriately conditioned sequence of learning rates $\eta_{t}$:
Algorithm 4 Stable (Practical) Implementation of Algorithm 2: SUBSAMPLE-Q: Learning
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Hyperparameter $k\in[n]$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$. Learning rate sequence $\{\eta_{t}\}_{t\in[T]}$ where $\eta_{t}\in(0,1]$.
1:  Let $\Delta=[k]$.
2:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{1},a_{g})=0$ for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}\times\mathcal{A}_{l}\times\mathcal{S}_{g}$.
3:  for $t=1$ to $T$ do
4:     for $(s_{g},s_{1},F_{z_{\Delta\setminus 1}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}^{(\mathcal{Z}_{l})}$ do
5:        for $(a_{g},a_{1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
6:           $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})\leftarrow(1-\eta_{t})\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})+\eta_{t}\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1})$
7:  Let the approximate policy be
$\widehat{\pi}_{k,m}^{T}(s_{g},s_{1},F_{s_{\Delta\setminus 1}})=\arg\max_{(a_{g},a_{1},a_{\Delta\setminus 1})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\Delta\setminus 1}},a_{g},a_{1}).$
Theorem B.16.
As $T\to\infty$, if $\sum_{t=1}^{T}\eta_{t}=\infty$, and $\sum_{t=1}^{T}\eta_{t}^{2}<\infty$, then $Q$-learning converges to the optimal $Q$ function asymptotically with probability $1$.
Furthermore, finite-time guarantees with the learning rate and sample complexity have been shown recently in Chen & Theja Maguluri (2022), which when adapted to our $\widehat{Q}_{k,m}$ framework in Algorithm 4 yields:
Theorem B.17 (Chen & Theja Maguluri (2022)).
For all $t\in[T]$ and $\epsilon>0$, if $\eta_{t}=(1-\gamma)^{4}\epsilon^{2}$ and $T=|\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}|\mathcal{S}_{g}||\mathcal{A}_{g}|/(1-\gamma)^{5}\epsilon^{2}$,
$\|\widehat{Q}_{k,m}^{T}-\widehat{Q}_{k,m}^{\mathrm{est}}\|\leq\epsilon.$
Definition B.18 (Star Graph $S_{n}$).
For $n\in\mathbb{N}$, the star graph $S_{n}$ is the complete bipartite graph $K_{1,n}$.
$S_{n}$ captures the graph density notion by saturating the set of neighbors of the central node. Such settings find applications beyond reinforcement learning as well (Chaudhari et al., 2024; Anand & Qu, 2024; Li et al., 2019). The cardinality of the search space simplex for the optimal policy is exponential in $n$, so it cannot be naively modeled by an MDP: we need to exploit the symmetry of the local agents. This intuition allows our subsampling algorithm to run in polylogarithmic time (in $n$). Some works leverage an exponential decaying property that truncates the search space for policies over immediate neighborhoods of agents; however, this still relies on the assumption that the graph neighborhood for the agent is sparse (Qu et al., 2020a, b); however, $S_{n}$ is not locally sparse; hence, previous methods do not apply to this problem instance.
[…]
Figure 3: Star graph $S_{n}$
Appendix C Proof Sketch
This section details an outline for the proof of Theorem 3.4, as well as some key ideas. At a high level, our SUBSAMPLE-MFQ framework recovers exact mean-field $Q$ learning and traditional value iteration when $k=n$ and as $m\to\infty$. Further, as $k\!\to\!n$, $\widehat{Q}_{k}^{*}$ should intuitively get closer to $Q^{*}$ from which the optimal policy is derived. Thus, the proof is divided into three major steps: firstly, we prove a Lipschitz continuity bound between $\widehat{Q}_{k}^{*}$ and $\widehat{Q}_{n}^{*}$ in terms of the total variation (TV) distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Next, we bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$. Finally, we bound the value differences between ${\pi}_{k}^{\mathrm{est}}$ and $\pi^{*}$ by bounding $Q^{*}(s,\pi^{*}(s))-Q^{*}(s,{\pi}_{k}^{\mathrm{est}}(s))$ and then using the performance difference lemma from Kakade & Langford (2002).
Step 1: Lipschitz Continuity Bound. To compare $\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g})$ with $Q^{*}(s,a_{g})$, we prove a Lipschitz continuity bound between $\widehat{Q}^{*}_{k}(s_{g},F_{s_{\Delta}},a_{g})$ and $\widehat{Q}^{*}_{k^{\prime}}(s_{g},F_{s_{\Delta^{\prime}}},a_{g})$ with respect to the TV distance measure between $s_{\Delta}\in\binom{s_{[n]}}{k}$ and $s_{\Delta^{\prime}}\in\binom{s_{[n]}}{k^{\prime}}$:
Theorem C.1 (Lipschitz continuity in $\widehat{Q}_{k}^{*}$).
For all $(s,a)\in\mathcal{S}\times\mathcal{A}$, $\Delta\in\binom{[n]}{k}$ and $\Delta^{\prime}\in\binom{[n]}{k^{\prime}}$,
$|\widehat{Q}^{*}_{k}(s_{g},F_{z_{\Delta}},a_{g})$ $-\widehat{Q}^{*}_{k^{\prime}}(s_{g},F_{z_{\Delta^{\prime}}},a_{g})|\leq{2(1-\gamma)^{-1}\|r_{l}(\cdot,\cdot)\|_{\infty}}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
We defer the proof of Theorem C.1 to Appendix D. See Fig. 4 for a comparison between the $\widehat{Q}_{k}^{*}$ learning and estimation process, and the exact ${Q}$-learning framework.
${\widehat{Q}^{0}_{k,m}(s_{g},F_{z_{\Delta}},a_{g})}$${Q^{*}(s_{g},s_{[n]},a_{g},a_{[n]})}$${\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g})}$${\widehat{Q}_{k}^{*}(s_{g},F_{z_{\Delta}},a_{g})}$${\widehat{Q}_{n}^{*}(s_{g},F_{z_{[n]}},a_{g})}$$\scriptstyle{\begin{subarray}{c}(1)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(2)\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(3)\\ \approx\end{subarray}}$$\scriptstyle{\begin{subarray}{c}(4)\\ =\end{subarray}}$
Figure 4: Flow of the algorithm and relevant analyses in learning $Q^{*}$. Here, (1) follows by performing Algorithm 2 (SUBSAMPLE-MFQ: Learning) on $\widehat{Q}_{k,m}^{0}$. (2) follows from Lemma 3.2. (3) follows from the Lipschitz continuity and total variation distance bounds in Theorems C.1 and C.2. Finally, (4) follows from noting that $\widehat{Q}_{n}^{*}=Q^{*}$.
Step 2: Bounding Total Variation (TV) Distance.
We bound the TV distance between $F_{z_{\Delta}}$ and $F_{z_{[n]}}$, where $\Delta\!\in\!\mathcal{U}\binom{[n]}{k}$. This task is equivalent to bounding the discrepancy between the empirical distribution and the distribution of the underlying finite population. When each $i\!\in\!\Delta$ is uniformly sampled without replacement, we use Lemma E.3 from Anand & Qu (2024) which generalizes the Dvoretzky-Kiefer-Wolfowitz (DKW) concentration inequality for empirical distribution functions. Using this, we show:
Theorem C.2.
Given a finite population $\mathcal{Z}=(z_{1},\dots,z_{n})$ for $\mathcal{Z}\in\mathcal{Z}_{l}^{n}$, let $\Delta\subseteq[n]$ be a uniformly random sample from $\mathcal{Z}$ of size $k$ chosen without replacement. Fix $\epsilon>0$. Then, for all $x\in\mathcal{Z}_{l}$:
$\Pr\bigg{[}\sup_{x\in\mathcal{Z}_{l}}\bigg{|}\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}{\{z_{i}=x\}}$ $-\frac{1}{n}\sum_{i\in[n]}\mathbbm{1}{\{z_{i}=x\}}\bigg{|}\leq\epsilon\bigg{]}\geq 1-2|\mathcal{Z}_{l}|e^{-\frac{2kn\epsilon^{2}}{n-k+1}}.$
Then, by Theorem C.2 and the definition of TV distance from Section 2, we have that for $\delta\in(0,1]$, with probability at least $1-\delta$,
$\mathrm{TV}(F_{s_{\Delta}},F_{s_{[n]}})\leq\sqrt{\frac{n-k+1}{8nk}\ln\frac{2|\mathcal{Z}_{l}|}{\delta}}$ (27)
We then apply this result to our MARL setting by studying the rate of decay of the objective function between the learned policy ${\pi}_{k}^{\mathrm{est}}$ and the optimal policy $\pi^{*}$ (Theorem 3.4).
Step 3: Performance Difference Lemma to Complete the Proof. As a consequence of the prior two steps and Lemma 3.2, $Q^{*}(s,a^{\prime})$ and $\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},F_{z_{\Delta}},a_{g}^{\prime})$ become similar as $k\to n$. We further prove that the value generated by the policies $\pi^{*}$ and ${\pi}_{k}^{\mathrm{est}}$ must also be very close (where the residue shrinks as $k\to n$). We then use the well-known performance difference lemma (Kakade & Langford, 2002) which we restate in Appendix F.1. A crucial theorem needed to use the performance difference lemma is a bound on $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\widehat{\pi}_{k}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))$. Therefore, we formulate and prove Theorem C.3 which yields a probabilistic bound on this difference, where the randomness is over the choice of $\Delta\in\binom{[n]}{k}$:
Theorem C.3.
For a fixed $s^{\prime}\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ and for $\delta\in(0,1]$, with probability atleast $1-2|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$:
$Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-Q^{*}(s^{\prime},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}^{\prime},F_{s_{\Delta}^{\prime}}))\leq\frac{2\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\sqrt{\frac{n-k+1}{2nk}\ln\left(\frac{2|\mathcal{Z}_{l}|}{\delta}\right)}+2\epsilon_{k,m}.$
We defer the proof of Theorem C.3 and finding optimal value of $\delta$ to Lemma F.7 in the Appendix. Using Theorem C.3 and the performance difference lemma leads to Theorem 3.4.
Appendix D Proof of Lipschitz-continuity Bound
This section proves a Lipschitz-continuity bound between $\widehat{Q}_{k}^{*}$ and $Q^{*}$ and includes a framework to compare $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}^{*}_{k}(s_{g},s_{\Delta},a_{g})$ and $Q^{*}(s,a_{g})$ in Lemma D.11.
Let $z_{i}:=(s_{i},a_{i})\in\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$ and $z_{\Delta}=\{z_{i}:i\in\Delta\}\in\mathcal{Z}_{l}^{k}$. For $z_{i}=(s_{i},a_{i})$, let $z_{i}(s)=s_{i}$, and $z_{i}(a)=a_{i}$. With abuse of notation, note that $\widehat{Q}_{k}^{T}(s_{g},a_{g},s_{\Delta},a_{\Delta})$ is equal to $\widehat{Q}_{k}^{T}(s_{g},a_{g},z_{\Delta})$.
The following definitions will be relevant to the proof of Theorem D.3.
Definition D.1 (Empirical Distribution Function).
For all $z\in\mathcal{Z}_{l}^{|\Delta|}$ and $s^{\prime},a^{\prime}\in\mathcal{S}_{l}\times\mathcal{A}_{l}$, where $\Delta\subseteq[n]$,
$F_{z_{\Delta}}(s^{\prime},a^{\prime})=\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}\{z_{i}(s)=s^{\prime},z_{i}(a)=a^{\prime}\}$
Definition D.2 (Total Variation Distance).
Let $P$ and $Q$ be discrete probability distribution over some domain $\Omega$. Then,
$\mathrm{TV}(P,Q)=\frac{1}{2}\left\|P-Q\right\|_{1}=\sup_{E\subseteq\Omega}\left|\Pr_{P}(E)-\Pr_{Q}(E)\right|$
Theorem D.3 ($\widehat{Q}_{k}^{T}$ is $\frac{2}{1-\gamma}\|r_{l}(\cdot,\cdot)\|_{\infty}$-Lipschitz continuous with respect to $F_{z_{\Delta}}$ in total variation distance).
Suppose $\Delta,\Delta^{\prime}\subseteq[n]$ such that $|\Delta|=k$ and $|\Delta^{\prime}|=k^{\prime}$. Then:
$\left|\widehat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}^{T}_{k^{\prime}}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
Proof.
We prove this inductively. First, note that $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ from the initialization step, which proves the lemma for $T=0$, since $\mathrm{TV}(\cdot,\cdot)\geq 0$. At $T=1$:
$|\widehat{Q}_{k}^{1}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$ $=\left|\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{0}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{\mathcal{T}}_{k^{\prime}}\widehat{Q}_{k^{\prime}}^{0}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|$
$=\bigg{|}r(s_{g},F_{s_{\Delta}},a_{g})+\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k}^{0}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a^{\prime}_{g})$
$\quad\quad\quad\quad-r(s_{g},F_{s_{\Delta^{\prime}}},a_{g})-\gamma\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k^{\prime}}^{0}(s^{\prime}_{g},F_{s^{\prime}_{\Delta^{\prime}}},a^{\prime}_{g})\bigg{|}$
$=|r(s_{g},a_{g},F_{z_{\Delta}})-r(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$
$=\left|\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{g},z_{i})-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{g},z_{i})\right|$
$=\bigg{|}\mathbb{E}_{z_{l}\sim F_{z_{\Delta}}}r_{l}(s_{g},z_{l})-\mathbb{E}_{z_{l}^{\prime}\sim F_{z_{\Delta^{\prime}}}}r_{l}(s_{g},z_{l}^{\prime})\bigg{|}$
In the first and second equalities, we use the time evolution property of $\widehat{Q}_{k}^{1}$ and $\widehat{Q}_{k^{\prime}}^{1}$ by applying the adapted Bellman operators $\widehat{\mathcal{T}}_{k}$ and $\widehat{\mathcal{T}}_{k^{\prime}}$ to $\widehat{Q}_{k}^{0}$ and $\widehat{Q}_{k^{\prime}}^{0}$, respectively, and expanding. In the third and fourth equalities, we note that $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$, and subtract the common ‘global component’ of the reward function.
Then, noting the general property that for any function $f:\mathcal{X}\to\mathcal{Y}$ for $|\mathcal{X}|<\infty$ we can write $f(x)=\sum_{y\in\mathcal{X}}f(y)\mathbbm{1}\{y=x\}$, we have:
$|\widehat{Q}^{1}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$ $=\left|\mathbb{E}_{z_{l}\sim F_{z_{\Delta}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}=z\}\right]-\mathbb{E}_{z_{l}^{\prime}\sim F_{z_{\Delta^{\prime}}}}\left[\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\mathbbm{1}\{z_{l}^{\prime}=z\}\right]\right|$
$=\bigg{|}\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\cdot(\mathbb{E}_{z_{l}\sim F_{z_{\Delta}}}\mathbbm{1}\{z_{l}=z\}-\mathbb{E}_{z_{l}^{\prime}\sim F_{z_{\Delta^{\prime}}}}\mathbbm{1}\{z_{l}^{\prime}=z\})\bigg{|}$
$=\bigg{|}\sum_{z\in\mathcal{Z}}r_{l}(s_{g},z)\cdot(F_{z_{\Delta}}(z)-F_{z_{\Delta^{\prime}}}(z))\bigg{|}$
$\leq\bigg{|}\max_{z\in\mathcal{Z}}r_{l}(s_{g},z)|\cdot\sum_{z\in\mathcal{Z}}|F_{z_{\Delta}}(z)-F_{z_{\Delta^{\prime}}}(z)\bigg{|}$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
The second equality follows from the linearity of expectations, and the third equality follows by noting that for any random variable $X\sim\mathcal{X}$, $\mathbb{E}_{X}\mathbbm{1}[X=x]=\Pr[X=x]$. The first inequality follows from an application of the triangle inequality and the Cauchy-Schwarz inequality, and the second inequality uses the definition of TV distance. Thus, at $T=1$, $\widehat{Q}$ is $(2\|r_{l}(\cdot,\cdot)\|_{\infty})$-Lipschitz continuous in TV distance, proving the base case.
Assume that for $T\leq t^{\prime}\in\mathbb{N}$:
$\left|\widehat{Q}^{T}_{k}(s_{g},a_{g},F_{z_{\Delta}})-\widehat{Q}_{k^{\prime}}^{T}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})\right|\leq\left(\sum_{t=0}^{T-1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
Then, inductively:
$|\widehat{Q}^{T+1}_{k}(s_{g},a_{g},F_{z_{\Delta}})$ $-\widehat{Q}_{k^{\prime}}^{T+1}(s_{g},a_{g},F_{z_{\Delta^{\prime}}})|$
$\leq\bigg{|}\frac{1}{|\Delta|}\sum_{i\in\Delta}r_{l}(s_{g},z_{i})-\frac{1}{|\Delta^{\prime}|}\sum_{i\in\Delta^{\prime}}r_{l}(s_{g},z_{i})\bigg{|}$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$\quad\quad\quad+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta})\sim\mathcal{J}_{k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta}^{\prime}})-\mathbb{E}_{(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}})\sim\mathcal{J}_{k^{\prime}}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},a_{g}^{\prime},F_{z_{\Delta^{\prime}}^{\prime}})\bigg{|}$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})+\gamma\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$=\left(\sum_{\tau=0}^{T}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
In the first inequality, we rewrite the expectations over the states as the expectation over the joint transition probabilities. The second inequality then follows from Lemma D.12. To apply it to Lemma D.12, we conflate the joint expectation over $(s_{g},s_{\Delta\cup\Delta^{\prime}})$ and reduce it back to the original form of its expectation. Finally, the third inequality follows from Lemma D.5.
By the inductive hypothesis, the claim is proven.∎
Definition D.4.
[Joint Stochastic Kernels] The joint stochastic kernel on $(s_{g},s_{\Delta})$ for $\Delta\subseteq[n]$ where $|\Delta|=k$ is defined as $\mathcal{J}_{k}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{l}^{k}\to[0,1]$, where
$\mathcal{J}_{k}(s_{g}^{\prime},s_{\Delta}^{\prime}|s_{g},a_{g},s_{\Delta},a_{\Delta}):=\Pr[(s_{g}^{\prime},s_{\Delta}^{\prime})|s_{g},a_{g},s_{\Delta},a_{\Delta}]$ (28)
Lemma D.5.
For all $T\in\mathbb{N}$, for any $a_{g}\in\mathcal{A}_{g},s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k},a_{\Delta}\in\mathcal{A}_{l}^{k},a^{\prime}_{\Delta}\in\mathcal{A}_{l}^{k}$, and for all joint stochastic kernels $\mathcal{J}_{k}$ as defined in Definition D.4:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta}})$ $-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime}_{g},a_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}})\bigg{|}$
$\quad\quad\quad\quad\quad\leq\left(\sum_{\tau=0}^{T-1}2\gamma^{\tau}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}\left(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}}\right)$
Proof.
We prove this inductively. At $T=0$, the statement is true since $\widehat{Q}_{k}^{0}(\cdot,\cdot,\cdot)=\widehat{Q}_{k^{\prime}}^{0}(\cdot,\cdot,\cdot)=0$ and $\mathrm{TV}(\cdot,\cdot)\geq 0$. At $T=1$,
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{1}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\bigg{[}r_{g}(s_{g}^{\prime},a_{g}^{\prime})+\frac{\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k}\bigg{]}$
$-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\bigg{[}r_{g}(s_{g}^{\prime},a_{g}^{\prime})+\frac{\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k^{\prime}}\bigg{]}\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\!\max_{a_{\Delta}^{\prime}}\frac{\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k}-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a^{\prime}_{\Delta^{\prime}}}\frac{\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})}{k^{\prime}}\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
In the last equality, we note that
$\max_{a_{\Delta}^{\prime}}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$ $=\sum_{i\in\Delta}\max_{a_{\Delta}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$
$=\sum_{i\in\Delta}\max_{a^{\prime}_{i}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$
$=\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})$
where $\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime}):=\max_{a_{i}^{\prime}}r_{l}(s_{i}^{\prime},a_{i}^{\prime},s_{g}^{\prime})$.
Then, analogously to an argument in Lemma B.3 of Anand & Qu (2024), we have:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\widetilde{r}_{l}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{x\in\mathcal{S}_{l}}\widetilde{r}_{l}(x,s_{g}^{\prime})F_{s^{\prime}_{\Delta}}(x)-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{x\in\mathcal{S}_{l}}\widetilde{r}_{l}(x,s_{g}^{\prime})F_{s^{\prime}_{\Delta^{\prime}}}(x)\bigg{|}$
$=\bigg{|}\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}}\sum_{x\in\mathcal{S}_{l}}\widetilde{r_{l}}(x,s_{g}^{\prime})\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}[F_{s^{\prime}_{\Delta}}(x)-F_{s^{\prime}_{\Delta^{\prime}}}(x)]\bigg{|}$
$\leq\|\widetilde{r_{l}}(\cdot,\cdot)\|_{\infty}\cdot\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}}\sum_{x\in\mathcal{S}_{l}}|\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}^{\prime}}(x)-\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}^{\prime}}(x)|$
$\leq\|{r_{l}}(\cdot,\cdot)\|_{\infty}\cdot\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}}\sum_{x\in\mathcal{S}_{l}}|\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}^{\prime}}(x)-\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}^{\prime}}(x)|$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}}\mathrm{TV}(\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}^{\prime}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}^{\prime}})$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z^{\prime}_{\Delta}})$
The first equality follows from noting that $f(x)=\sum_{x^{\prime}\in\mathcal{X}}f(x^{\prime})\mathbbm{1}\{x=x^{\prime}\}$ and from Fubini-Tonelli’s inequality which allows us to swap the order of summations as the summand is finite. The second equality uses the law of total expectation. The first inequality uses Jensen’s inequality and the triangle inequality. The second inequality uses $\|\widetilde{r}_{l}(\cdot,\cdot)\|_{\infty}\leq\|r_{l}(\cdot,\cdot)\|_{\infty}$ which holds as $\|{r}_{l}\|_{\infty}$ is the infinite-norm of the local reward functions and is therefore atleast as large any other element in the image of $r_{l}$. The third inequality follows from the definition of total variation distance, and the final inequality follows from Lemma D.7. This proves the base case.
Then, assume that for $T\leq t^{\prime}\in\mathbb{N}$, for all joint stochastic kernels $\mathcal{J}_{k}$ and $\mathcal{J}_{k^{\prime}}$, and for all $a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}$:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{T}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})$ $-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
$\leq 2\left(\sum_{t=0}^{T-1}\!\gamma^{t}\right)\!\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
For the remainder of the proof, we adopt the shorthand $\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}}}$ to denote $\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{|\Delta|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}$, and $\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim{\mathcal{J}}}$ to denote $\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{|\Delta|}(\cdot,\cdot|s_{g}^{\prime},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})}$. Then, inductively, we have:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{T+1}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{T+1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\underbrace{\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\bigg{[}r_{\Delta}(s_{g}^{\prime},a_{g}^{\prime},s^{\prime}_{\Delta},a^{\prime}_{\Delta})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\bigg{]}}_{\text{(I)}}$
$-\underbrace{\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime}}\bigg{[}r_{\Delta^{\prime}}(s_{g}^{\prime},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},a_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{\Delta^{\prime}}^{\prime})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime})\bigg{]}}_{\text{(II)}}\bigg{|}$
Let
$\widetilde{a}_{g},\widetilde{a}_{\Delta}=\arg\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\bigg{[}r_{\Delta}(s_{g}^{\prime},a_{g}^{\prime},s^{\prime}_{\Delta},a^{\prime}_{\Delta})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\bigg{]},$
and
$\widehat{a}_{g},\widehat{a}_{\Delta^{\prime}}=\arg\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k}}\bigg{[}r_{\Delta^{\prime}}(s_{g}^{\prime},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{\Delta^{\prime}}^{\prime})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime})\bigg{]}$
Then, define $\widetilde{a}_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}$ by
$\widetilde{a}_{i}=\begin{cases}\widetilde{a}_{i},&\text{if $i\in\Delta$},\\ \widehat{a}_{i},&\text{if $i\in\Delta^{\prime}\setminus\Delta$}\end{cases}$
Similarly, define $\widehat{a}_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}$ by
$\widehat{a}_{i}=\begin{cases}\widehat{a}_{i},&i\in\Delta^{\prime}\\ \widetilde{a}_{i},&i\in\Delta\setminus\Delta^{\prime}\end{cases}$
Suppose (I)$\geq$(II). Then,
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{T+1}(s^{\prime}_{g},a_{g}^{\prime},s_{\Delta}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}}}\widehat{Q}_{k^{\prime}}^{T+1}(s^{\prime}_{g},a_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a^{\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\widehat{Q}_{k}^{T+1}(s^{\prime}_{g},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\widehat{Q}_{k^{\prime}}^{T+1}(s^{\prime}_{g},\widehat{a}_{g},s^{\prime}_{\Delta^{\prime}},\widehat{a}_{\Delta^{\prime}})\bigg{|}$
$\leq\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\widehat{Q}_{k}^{T+1}(s^{\prime}_{g},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\widehat{Q}_{k^{\prime}}^{T+1}(s^{\prime}_{g},\widetilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\bigg{[}r_{\Delta}(s_{g}^{\prime},s_{\Delta}^{\prime},\widetilde{a}_{g},\widetilde{a}_{\Delta})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},a_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})\bigg{]}$
$-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\bigg{[}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{g},\widetilde{a}_{\Delta^{\prime}})+\gamma\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},s{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
$\leq\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}r_{\Delta}(s_{g}^{\prime},s_{\Delta}^{\prime},\widetilde{a}_{g},\widetilde{a}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}r_{\Delta^{\prime}}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{g},\widetilde{a}_{\Delta^{\prime}})\bigg{|}$
$+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s_{\Delta}^{\prime},\widetilde{a}_{\Delta})}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})$
$-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime}}^{\prime\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g}^{\prime},\widetilde{a}_{g},s^{\prime}_{\Delta^{\prime}},\widetilde{a}_{\Delta^{\prime}})}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
$=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim{\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim{\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})\bigg{|}$
$+\gamma\bigg{|}\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim{\widetilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}}\max_{a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime}}\widehat{Q}_{k}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s_{\Delta}^{\prime\prime},{a}^{\prime\prime}_{\Delta})-\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}^{\prime})\sim{\widetilde{\mathcal{J}}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}}\max_{a_{g}^{\prime\prime},a_{\Delta^{\prime}}^{\prime\prime}}\widehat{Q}_{k^{\prime}}^{T}(s^{\prime\prime}_{g},{a}^{\prime\prime}_{g},s^{\prime\prime}_{\Delta^{\prime}},{a}^{\prime\prime}_{\Delta^{\prime}})\bigg{|}$
$\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})+\gamma\left(\sum_{t=0}^{T}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
$=\left(\sum_{t=0}^{T+1}2\gamma^{t}\right)\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
The first equality rewrites the equations with their respective maximizing actions. The first inequality upper-bounds this difference by allowing all terms to share the common action $\widetilde{a}$. Using the Bellman equation, the second equality expands $\widehat{Q}_{k}^{T+1}$ and $\widehat{Q}_{k^{\prime}}^{T+1}$. The second inequality follows from the triangle inequality. The third equality follows by subtracting the common $r_{g}(s_{g}^{\prime},\widetilde{a_{g}})$ terms from the reward and noting that the two expectation terms $\mathbb{E}_{s_{g}^{\prime},s_{\Delta}^{\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta^{\prime\prime}}\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},\widetilde{a}_{g},s_{\Delta},\widetilde{a}_{\Delta})}$ can be combined into a single expectation $\mathbb{E}_{s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime}\sim\widetilde{\mathcal{J}}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}$.
In the special case where $a_{\Delta}=\widetilde{a}_{\Delta}$, we derive a closed form expression for $\widetilde{\mathcal{J}}_{k}$ in Lemma D.10. To justify the third inequality, the second term follows from the induction hypothesis and the first term follows from the following derivation:
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})\bigg{|}$
$\quad\quad\quad=\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\frac{1}{k}\sum_{i\in\Delta}r^{\widetilde{a}}_{l}(s_{i}^{\prime},s_{g}^{\prime})-\mathbb{E}_{(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{k^{\prime}}(\cdot,\cdot|s_{g},a_{g},s_{\Delta^{\prime}},a_{\Delta^{\prime}})}\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}r_{l}^{\widetilde{a}}(s_{i}^{\prime},s_{g}^{\prime})\bigg{|}$
$\quad\quad\quad\leq\|r_{l}^{\widetilde{a}}(\cdot,\cdot)\|_{\infty}\cdot\mathbb{E}_{s_{g}^{\prime}\sim\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathrm{TV}(\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}}F_{s_{\Delta^{\prime}}})$
$\quad\quad\quad\leq 2\|r_{l}(\cdot,\cdot)\|_{\infty}\cdot\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
The above derivation follows from the same argument as in the base case where $r_{l}^{\widetilde{a}}(s_{i}^{\prime},s_{g}^{\prime}):=r_{l}(s_{i}^{\prime},s_{g}^{\prime},\widetilde{a}_{i})$ for any $i\in\Delta\cup\Delta^{\prime}$. Similarly, if (I)$<$(II), an analogous argument that replaces $\widetilde{a}_{\Delta}$ with $\widehat{a}_{\Delta}$ yields the same result.
Therefore, by the induction hypothesis, the claim is proven.∎
Remark D.6.
Given a joint transition probability function $\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}$ as defined in Definition D.4, we can recover the transition function for a single agent $i\in\Delta\cup\Delta^{\prime}$ given by $\mathcal{J}_{1}$ using the law of total probability and the conditional independence between $s_{i}$ and $s_{g}\cup s_{[n]\setminus i}$ in Eq. 29. This characterization is crucial in Lemma D.7 and Lemma D.5:
$\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})=\sum_{s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}\sim\mathcal{S}_{l}^{|\Delta\cup\Delta^{\prime}|-1}}\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}\circ s^{\prime}_{i}|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})$ (29)
Here, we use a conditional independence property proven in Lemma D.9.
Lemma D.7.
$\mathrm{TV}\bigg{(}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}\bigg{)}\leq\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
Proof.
From the definition of total variation distance, we have:
$\mathrm{TV}\bigg{(}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}}}F_{s^{\prime}_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}}}F_{s^{\prime}_{\Delta^{\prime}}}\bigg{)}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\bigg{|}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)-\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta^{\prime}}}(x)\bigg{|}$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\right|$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})\mathbbm{1}\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}\right|$
$\quad\quad=\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right|$
The second equality uses Lemma D.8. The third equality uses the property that each local agent can only have one state/action pair, and the fourth equality vectorizes the indicator random variables. Then,
$\mathrm{TV}\bigg{(}\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}}}F_{s^{\prime}_{\Delta}},\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}}}F_{s^{\prime}_{\Delta^{\prime}}}\bigg{)}$
$\quad\quad\quad\quad\leq\frac{1}{2}\sum_{x\in\mathcal{S}_{l}}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right|$
$\quad\quad\quad\quad=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}$
$\quad\quad\quad\quad\leq\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\vec{\mathbbm{1}}_{\left\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\right\}}\right\|_{1}\cdot\|\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},\cdot,\cdot)\|_{1}$
$\quad\quad\quad\quad=\frac{1}{2}\sum_{a_{l}\in\mathcal{A}_{l}}\sum_{s_{l}\in\mathcal{S}_{l}}\left\|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right\|_{1}$
$\quad\quad\quad\quad=\frac{1}{2}\sum_{(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}-\frac{1}{k^{\prime}}\sum_{i\in\Delta^{\prime}}\mathbbm{1}\{\begin{subarray}{c}a_{i}=a_{l}\\ s_{i}=s_{l}\end{subarray}\}\right|$
$\quad\quad\quad\quad=\frac{1}{2}\sum_{z_{l}\in\mathcal{S}_{l}\times\mathcal{A}_{l}}\left|F_{z_{\Delta}}(z_{l})-F_{z_{\Delta^{\prime}}}(z_{l})\right|$
$\quad\quad\quad\quad:=\mathrm{TV}(F_{z_{\Delta}},F_{z_{\Delta^{\prime}}})$
The first inequality uses the triangle inequality. The second inequality and fifth equality follow from Hölder’s inequality and the sum of the probabilities from the stochastic transition function $\mathcal{J}_{1}$ being equal to $1$. The sixth equality uses Fubini-Tonelli’s theorem which applies as the total variation distance measure is bounded from above by $1$. The final equality recovers the total variation distance for the variable $z=(s,a)\in\mathcal{S}_{l}\times\mathcal{A}_{l}$ across agents $\Delta$ and $\Delta^{\prime}$.∎
Lemma D.8.
$\mathbb{E}_{s^{\prime}_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)=\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})$
Proof.
By expanding on the definition of the empirical distribution function $F_{s_{\Delta^{\prime}}}(x)=\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}\{s_{i}=x\}$, we have:
$\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}F_{s^{\prime}_{\Delta}}(x)$ $=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$=\frac{1}{k}\sum_{i\in\Delta}\mathbb{E}_{s_{\Delta\cup\Delta^{\prime}}|s_{g}^{\prime}\sim\mathcal{J}_{1}(\cdot|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})}\mathbbm{1}\{s_{i}^{\prime}=x\}$
$=\frac{1}{k}\sum_{i\in\Delta}\mathcal{J}_{1}(x|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i})$
The second equality follows from the conditional independence of $s_{i}^{\prime}$ from $s_{\Delta\cup\Delta^{\prime}\setminus i}$ from Lemma D.9, and the final equality uses the fact that the expectation of an indicator random variable is the probability distribution function of the random variable. ∎
Lemma D.9.
The distribution $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}|s_{g}^{\prime},s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}}$ is conditionally independent to the distribution $s_{i}^{\prime}|s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i}$ for any $i\in\Delta\cup\Delta^{\prime}$.
Proof.
We direct the interested reader to the Bayes-Ball theorem in Shachter (2013) for proving conditional independence. For ease of exposition, we restate the two rules in Shachter (2013) that introduce the notion of $d$-separations which implies conditional independence. Suppose we have a causal graph $G=(V,E)$ where the vertex set $V=[p]$ for $p\in\mathbb{N}$ is a set of variables and the edge set $E\subseteq 2^{V}$ denotes dependence through connectivity. Then, the two rules that establish $d$-separations are as follows:
  1. 1.
    For $x,y\in V$, we say that $x,y$ are $d$-connected if there exists a path $(x,\dots,y)$ that can be traced without traversing a pair of arrows that point at the same vertex.
  2. 2.
    We say that $x,y\in V$ are $d$-connected, conditioned on a set of variables $Z\subseteq V$, if there is a path $(x,\dots,y)$ that does not contain an event $z\in Z$ that can be traced without traversing a pair of arrows that point at the same vertex.
If $x,y\in V$ is not $d$-connected through any such path, then $x,y$ is $d$-separated, which implies conditional independence.
Let $Z=\{s_{g}^{\prime},s_{g},a_{g},s_{i},a_{i}\}$ be the set of variables we condition on. Then, the below figure demonstrates the causal graph for the events of interest.
$a_{g}$$s_{\Delta\cup\Delta^{\prime}\setminus i}$$s_{g}$$s_{i}$$a_{\Delta\cup\Delta^{\prime}\setminus i}$$a_{i}$$a_{g}^{\prime}$$s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$$s^{\prime}_{g}$$s^{\prime}_{i}$$a^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$$a^{\prime}_{i}$
Figure 5: Causal graph to demonstrate the dependencies between variables.
  1. 1.
    Observe that all paths (through undirected edges) stemming from $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}\to s_{\Delta\cup\Delta^{\prime}\setminus i}$ pass through $s_{i}\in Z$ which is blocked.
  2. 2.
    All other paths from $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ to $s_{i}^{\prime}$ pass through $s_{g}\cup s_{g}^{\prime}\in Z$.
Therefore, $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are $d$-separated by $Z$. Hence, by Shachter (2013), $s^{\prime}_{\Delta\cup\Delta^{\prime}\setminus i}$ and $s_{i}^{\prime}$ are conditionally independent.∎
Lemma D.10.
For any joint transition probability function $\mathcal{J}_{k}$ on $s_{g}\in\mathcal{S}_{g},s_{\Delta}\in\mathcal{S}_{l}^{k}$, $a_{\Delta}\in\mathcal{A}_{l}^{k}$ where $|\Delta|=k$, given by $\mathcal{J}_{k}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}\times\mathcal{S}_{g}\times\mathcal{A}_{g}\times\mathcal{S}_{l}^{|\Delta|}\times\mathcal{A}_{l}^{|\Delta|}\to[0,1]$, we have:
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$\quad\quad\quad=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
Proof.
We start by expanding the expectations:
$\mathbb{E}_{(s_{g}^{\prime},s_{\Delta}^{\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\left[\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}(\cdot,\cdot|s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})\right]$
$\quad\quad\quad=\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]\max_{\begin{subarray}{c}a_{g}^{\prime\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$\quad\quad\quad=\sum_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}^{2}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}^{T}_{k}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
$\quad\quad\quad=\mathbb{E}_{(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})\sim\mathcal{J}_{k}^{2}(\cdot,\cdot|s_{g},a_{g},s_{\Delta},a_{\Delta})}\max_{a_{g}^{\prime\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},a_{g}^{\prime\prime},a_{\Delta}^{\prime\prime})$
In the second equality, we note that the right-stochasticity of $\mathcal{J}_{k}$ implies the right-stochasticity of $\mathcal{J}_{k}^{2}$. Further, observe that $\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{\Delta}^{\prime},a_{\Delta}]$ denotes the probability of the transitions $(s_{g},s_{\Delta})\to(s_{g}^{\prime},s_{\Delta}^{\prime})\to(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ with actions $a_{g},a_{\Delta}$ at each step, where the joint state evolution is governed by $\mathcal{J}_{k}$. Thus,
$\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{g}^{\prime},a_{\Delta}]=\mathcal{J}_{k}^{2}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]$
since $\sum_{(s_{g}^{\prime},s_{\Delta}^{\prime})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{|\Delta|}}\mathcal{J}_{k}[s_{g}^{\prime},s_{\Delta}^{\prime},s_{g},a_{g},s_{\Delta},a_{\Delta}]\mathcal{J}_{k}[s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime},s_{g}^{\prime},a_{g},s_{g}^{\prime},a_{\Delta}]$ is the stochastic probability function corresponding to the two-step evolution of the joint states from $(s_{g},s_{\Delta})$ to $(s_{g}^{\prime\prime},s_{\Delta}^{\prime\prime})$ under actions $a_{g},a_{\Delta}$. This can be thought of as an analogously to the fact that a $1$ on the $(i,j)$’th entry on the square of a $0/1$ adjacency matrix of a graph represents the fact that there is a path of distance $2$ between vertices $i$ and $j$. Finally, the third equality recovers the definition of the expectation, with respect to the joint probability function $\mathcal{J}_{k}^{2}$. ∎
We next show that $\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})=\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=Q^{*}(s,a)$.
Lemma D.11.
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
Proof.
We bound the differences between $\widehat{Q}_{k}^{T}$ at each Bellman iteration of our approximation to $Q^{*}$.
Note that:
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
$=\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
$=r_{[n]}(s_{g},s_{[n]},a_{g})+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim P_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n])\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{[n]}^{\prime}\in\mathcal{A}_{l}^{n}}Q^{*}(s^{\prime},a^{\prime})$
$\quad\quad\quad\quad-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\bigg{[}r_{\Delta}(s_{g},s_{\Delta},a_{g},a_{\Delta})+\gamma\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i}s_{g}),\forall i\in\Delta\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}Q_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
Next, observe that $r_{[n]}(s_{g},s_{[n]},a_{g},a_{[n]})=\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}r_{[\Delta]}(s_{g},s_{\Delta},a_{g},a_{\Delta})$. To prove this, we write:
$\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}r_{[\Delta]}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ $=\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}(r_{g}(s_{g},a_{g})+\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i},a_{i},s_{g}))$
$=r_{g}(s_{g},a_{g})+\frac{{n-1\choose k-1}}{k{n\choose k}}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
$=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g}):=r_{[n]}(s_{g},s_{[n]},a_{g},a_{[n]})$
[…]
So, we have that:
$\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{[n]},a_{g},a_{\Delta})\bigg{]}$
$=\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathcal{T}Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{T-1}(s_{g},s_{\Delta},a_{g},a_{\Delta})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\bigg{[}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in\Delta\end{subarray}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$=\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\bigg{[}\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathbb{E}_{\begin{subarray}{c}s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g}),\\ s_{i}^{\prime}\sim P_{l}(\cdot|s_{i},a_{i},s_{g}),\forall i\in[n]\end{subarray}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{[n]}^{\prime}\in\mathcal{A}_{l}^{n}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T-1}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})\bigg{]}$
$\leq\gamma\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T-1}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
We justify the first inequality by noting the general property that for positive vectors $v,v^{\prime}$ for which $v\succeq v^{\prime}$ which follows from the triangle inequality:
$\bigg{\|}v-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}v^{\prime}\bigg{\|}_{\infty}$ $\geq\bigg{|}\|v\|_{\infty}-\bigg{\|}\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}v^{\prime}\bigg{\|}_{\infty}\bigg{|}$
$=\|v\|_{\infty}-\bigg{\|}\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}v^{\prime}\bigg{\|}_{\infty}$
$\geq\|v\|_{\infty}-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\|v^{\prime}\|_{\infty}$
Thus, applying this bound recursively, we get:
$Q^{*}(s,a)-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$ $\leq\gamma^{T}\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}\bigg{[}Q^{*}(s^{\prime},a^{\prime})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{0}(s^{\prime}_{g},s^{\prime}_{\Delta},a^{\prime}_{g},a^{\prime}_{\Delta})\bigg{]}$
$=\gamma^{T}\sup_{(s^{\prime},a^{\prime})\in\mathcal{S}\times\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})$
$=\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
The first inequality follows from the $\gamma$-contraction property of the update procedure, and the ensuing equality follows from our bound on the maximum possible value of $Q$ from Lemma B.5 and noting that $\widehat{Q}_{k}^{0}:=0$. Therefore, as $T\to\infty$,
$Q^{*}(s,a_{g})-\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g})\to 0,$
which proves the lemma.∎
Corollary D.11.1.
Since $\frac{1}{{n\choose k}}\sum_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta}):=\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, we therefore get:
$Q^{*}(s,a)-\lim_{T\to\infty}\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq\lim_{T\to\infty}\gamma^{T}\cdot\frac{\widetilde{r}}{1-\gamma}$
Consequently,
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq 0$
Further, we have that
$Q^{*}(s,a)-\mathbb{E}_{\Delta\in{[n]\choose k}}\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0,$
since $\widehat{Q}_{k}^{*}(s_{g},s_{\Delta},a_{g},a_{\Delta})\leq Q^{*}(s,a)$.
Lemma D.12.
The absolute difference between the expected maximums between $\widehat{Q}_{k}$ and $\widehat{Q}_{k^{\prime}}$ is atmost the maximum of the absolute difference between $\widehat{Q}_{k}$ and $\widehat{Q}_{k^{\prime}}$, where the expectations are taken over any joint distributions of states $\mathcal{J}$, and the maximums are taken over the actions.
$\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$ $-\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a^{\prime}_{\Delta^{\prime}})\bigg{]}\bigg{|}$
$\leq\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a^{\prime}_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}\bigg{[}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{\prime},a_{\Delta}^{\prime})$ $-\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{]}\bigg{|}$
Proof.
Denote:
$a_{g}^{*},a_{\Delta}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta}},a_{g}^{\prime},a_{\Delta}^{\prime})$
$\widetilde{a}_{g}^{*},\widetilde{a}_{\Delta^{\prime}}^{*}:=\arg\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}\end{subarray}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{s^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})$
We extend $a_{\Delta}^{*}$ to $a_{\Delta^{\prime}}^{*}$ by letting $a_{\Delta^{\prime}\setminus\Delta}^{*}$ be the corresponding $\Delta\setminus\Delta^{\prime}$ variables from $\widetilde{a}_{\Delta^{\prime}}^{*}$ in $\mathcal{A}_{l}^{|\Delta^{\prime}\setminus\Delta|}$. For the remainder of this proof, we adopt the shorthand $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}$ to refer to $\mathbb{E}_{(s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime})\sim\mathcal{J}_{|\Delta\cup\Delta^{\prime}|}(\cdot,\cdot|s_{g},a_{g},s_{\Delta\cup\Delta^{\prime}},a_{\Delta\cup\Delta^{\prime}})}$. Then, if $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})>0$, we have:
$\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta}},a_{g}^{\prime})$ $-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},F_{z^{\prime}_{\Delta^{\prime}}},a_{g}^{\prime})\bigg{|}$
$=\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s_{\Delta}^{\prime},a_{g}^{*},a^{*}_{\Delta})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},\widetilde{a}_{g}^{*},\widetilde{a}^{*}_{\Delta^{\prime}})$
$\leq\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}^{T}_{k}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{*},a_{\Delta}^{*})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{*},a_{\Delta^{\prime}}^{*})$
$\leq\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},\\ a_{\Delta\cup\Delta^{\prime}}\in\mathcal{A}_{l}^{|\Delta\cup\Delta^{\prime}|}\end{subarray}}\bigg{|}\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta^{\prime}},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})\bigg{|}$
Similarly, if $\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k}^{T}(s_{g}^{\prime},s^{\prime}_{\Delta},a_{g}^{\prime},a_{\Delta}^{\prime})-\mathbb{E}_{s_{g}^{\prime},s_{\Delta\cup\Delta^{\prime}}^{\prime}}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta^{\prime}}^{\prime}\in\mathcal{A}_{l}^{k^{\prime}}}\widehat{Q}_{k^{\prime}}^{T}(s_{g}^{\prime},s_{\Delta^{\prime}}^{\prime},a_{g}^{\prime},a_{\Delta^{\prime}}^{\prime})<0$, an analogous argument by replacing $a_{g}^{*}$ with $\widetilde{a}_{g}^{*}$ and $a_{\Delta}^{*}$ with $\widetilde{a}_{\Delta}^{*}$ yields an identical bound. ∎
Lemma D.13.
Suppose $z,z^{\prime}\geq 1$. Consider functions $\Gamma:\Theta_{1}\times\Theta_{2}\times\dots\times\Theta_{z}\times\Theta^{*}\to\mathbb{R}$ and $\Gamma^{\prime}:\Theta_{1}^{\prime}\times\Theta_{2}^{\prime}\times\dots\times\Theta^{\prime}_{z^{\prime}}\times\Theta^{*}\to\mathbb{R}$, where $\Theta_{1},\dots,\Theta_{z}$ and $\Theta_{1}^{\prime},\dots,\Theta^{\prime}_{z^{\prime}}$ are finite sets. Consider a probability distribution function $\mu_{\Theta_{i}}$ for $i\in[z]$ and $\mu^{\prime}_{\Theta_{i}}$ for $i\in[z^{\prime}]$. Then:
$\bigg{|}\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})$ $-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})\bigg{|}$
$\leq\max_{\theta^{*}\in\Theta^{*}}\bigg{|}\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})\bigg{|}$
Proof.
Let $\widehat{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})$ and $\widetilde{\theta}^{*}:=\arg\max_{\theta^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})$.
If $\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}},\dots,\theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}},\dots,\theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})>0$, then:
$\bigg{|}\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})$ $-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})\bigg{|}$
$=\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\widehat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\widetilde{\theta}^{*})$
$\leq\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},\widehat{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\widehat{\theta}^{*})$
$\leq\max_{\theta^{*}\in\Theta^{*}}\bigg{|}\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}}\\ \dots\\ \theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\Gamma(\theta_{1},\dots,\theta_{z},{\theta}^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}}\\ \dots\\ \theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},{\theta}^{*})\bigg{|}$
Here, we replace each $\theta^{*}$ with the maximizers of their corresponding terms, and upper bound them by the maximizer of the larger term. Next, we replace $\widehat{\theta^{*}}$ in both expressions with the maximizer choice $\theta^{*}$ from $\Theta^{*}$, and further bound the expression by its absolute value.
If $\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu_{\Theta_{1}},\dots,\theta_{z}\sim\mu_{\Theta_{z}}\end{subarray}}\max_{\theta^{*}\in\Theta^{*}}\Gamma(\theta_{1},\dots,\theta_{z},\theta^{*})-\mathbb{E}_{\begin{subarray}{c}\theta_{1}\sim\mu^{\prime}_{\Theta^{\prime}_{1}},\dots,\theta_{z^{\prime}}\sim\mu^{\prime}_{\Theta^{\prime}_{z^{\prime}}}\end{subarray}}\max_{{\theta}^{*}\in\Theta^{*}}\Gamma^{\prime}(\theta_{1},\dots,\theta_{z^{\prime}},\theta^{*})<0$, then an analogous argument that replaces $\widehat{\theta}^{*}$ with $\widetilde{\theta}^{*}$ yields the same result. ∎
Appendix E Bounding Total Variation Distance
As $|\Delta|\to n$, we prove that the total variation (TV) distance between the empirical distribution of $z_{\Delta}$ and $z_{[n]}$ goes to 00. Here, recall that $z_{i}\in\mathcal{Z}=\mathcal{S}_{l}\times\mathcal{A}_{l}$, and $z_{\Delta}=\{z_{i}:i\in\Delta\}$ for $\Delta\in{[n]\choose k}$. Before bounding the total variation distance between $F_{z_{\Delta}}$ and $F_{z_{\Delta^{\prime}}}$, we first introduce Lemma C.5 of Anand & Qu (2024) which can be viewed as a generalization of the Dvoretzky-Kiefer-Wolfowitz concentration inequality, for sampling without replacement. We first make an important remark.
Remark E.1.
First, observe that if $\Delta$ is an independent random variable uniformly supported on (${[n]\choose k}$, then $s_{\Delta}$ and $a_{\Delta}$ are also independent random variables uniformly supported on the global state (${s_{[n]}\choose k}$ and the global action (${a_{[n]}\choose k}$. To see this, let $\psi_{1}:[n]\to\mathcal{S}_{l}$ where $\psi_{1}(i)=s_{i}$ and $\xi_{1}:[n]\to\mathcal{A}_{l}$ where $\xi_{1}(i)=a_{i}$. This naturally extends to $\psi_{k}:[n]^{k}\to\mathcal{S}_{l}^{k}$, where $\psi_{k}(i_{1},\dots,i_{k})=(s_{i_{1}},\dots,s_{i_{k}})$ and $\xi_{k}:[n]^{k}\to\mathcal{A}_{l}^{k}$, where $\xi_{k}(i_{1},\dots,i_{k})=(a_{i_{1}},\dots,a_{i_{k}})$ for all $k\in[n]$. Then, the independence of $\Delta$ implies the independence of the generated $\sigma$-algebra. Further, $\psi_{k}$ and $\xi_{k}$ (which are a Lebesgue measurable function of a $\sigma$-algebra) are sub-algebras, implying that $s_{\Delta}$ and $a_{\Delta}$ must also be independent random variables.
For reference, we present the multidimensional Dvoretzky-Kiefer-Wolfowitz (DKW) inequality (Dvoretzky et al. (1956); Massart (1990); Naaman (2021)) which bounds the difference between an empirical distribution function for a set $B_{\Delta}$ and $B_{[n]}$ when each element of $\Delta$ for $|\Delta|=k$ is sampled uniformly at random from $[n]$ with replacement.
Theorem E.2 (Multi-dimensional Dvoretzky-Kiefer-Wolfowitz (DFW) inequality (Dvoretzky et al., 1956; Naaman, 2021)).
Suppose $B\subset\mathbb{R}^{d}$ and $\epsilon>0$. If $\Delta\subseteq[n]$ is sampled uniformly with replacement, then
$\Pr\left[\sup_{x\in B}\left|\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}\{B_{i}=x\}-\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\{B_{i}=x\}\right|<\epsilon\right]\geq 1-d(n+1)e^{-2|\Delta|\epsilon^{2}}\cdot$
Lemma C.5 of Anand & Qu (2024) generalizes the DKW inequality for sampling without replacement:
Lemma E.3 (Sampling without replacement analogue of the DKW inequality, Lemma C.5 in Anand & Qu (2024)).
Consider a finite population $\mathcal{X}=(x_{1},\dots,x_{n})\in\mathcal{B}_{l}^{n}$ where $\mathcal{B}_{l}$ is a finite set. Let $\Delta\subseteq[n]$ be a random sample of size $k$ chosen uniformly and without replacement. Then, for all $x\in\mathcal{B}_{l}$:
$\Pr\left[\sup_{x\in\mathcal{B}_{l}}\left|\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}{\{x_{i}=x\}}-\frac{1}{n}\sum_{i\in[n]}\mathbbm{1}{\{x_{i}=x\}}\right|<\epsilon\right]\geq 1-2|\mathcal{B}_{l}|e^{-\frac{2|\Delta|n\epsilon^{2}}{n-|\Delta|+1}}$
Lemma E.4.
With probability atleast $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$, $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})\leq\epsilon$.
Proof.
Recall that $z:=(s_{l},a_{l})\in\mathcal{S}_{l}\times\mathcal{A}_{l}$. From Lemma E.3, substituting $\mathcal{B}_{l}=\mathcal{Z}_{l}$ yields:
$\Pr\left[\sup_{z_{l}\in\mathcal{Z}_{l}}\left|F_{z_{\Delta}}(z_{l})-F_{z_{[n]}}(z_{l})\right|\leq 2\epsilon\right]$ $\geq 1-2|\mathcal{Z}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$
$=1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}},$
which yields the proof.∎
We now present an alternate bound for the total variation distance, where the distance actually goes to 00 as $|\Delta|\to n$. For this, we use the fact that the total variation distance between two product distributions is subadditive.
Lemma E.5 (Lemma B.8.1 of Ghosal & van der Vaart (2017). Subadditivity of TV distance for Product Distributions).
Let $P$ and $Q$ be product distributions over some domain $S$. Let $\alpha_{1},\dots,\alpha_{d}$ be the marginal distributions of $P$ and $\beta_{1},\dots,\beta_{q}$ be the marginal distributions of $Q$. Then,
$\|P-Q\|_{1}\leq\sum_{i=1}^{d}\|\alpha_{i}-\beta_{i}\|_{1}$
Furthermore, we will need the following lemma.
Lemma E.6 (Data Processing Inequality.).
Let $A$ and $B$ be random variables over some domain $S$. Let $f$ be some function (not necessarily deterministically) mapping from $S$ to any codomain $T$. Then,
$\mathrm{TV}(f(A),f(B))\leq\mathrm{TV}(A,B)$
Lemma E.7.
$\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})\leq\sqrt{1-\frac{|\Delta|}{n}}$
Proof.
By the symmetry of the total variation distance metric, we have that
$\mathrm{TV}(F_{z_{[n]}},F_{z_{\Delta}})=\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}}).$
From the Bretagnolle-Huber inequality Tsybakov (2008) we have that
$\mathrm{TV}(f,g)=\sqrt{1-e^{-D_{\mathrm{KL}}(f\|g)}}.$
Here, $D_{\mathrm{KL}}(f\|g)$ is the Kullback-Leibler (KL) divergence metric between probability distributions $f$ and $g$ over the sample space, which we denote by $\mathcal{X}$ and is given by
$D_{\mathrm{KL}}(f\|g):=\sum_{x\in\mathcal{X}}f(x)\ln\frac{f(x)}{g(x)}$ (30)
Thus, from Eq. 30:
$D_{\mathrm{KL}}(F_{z_{\Delta}}\|F_{z_{[n]}})$ $=\sum_{z\in\mathcal{Z}_{l}}\left(\frac{1}{|\Delta|}\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\right)\ln\frac{n\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}}{|\Delta|\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}}$
$=\frac{1}{|\Delta|}\sum_{z\in\mathcal{Z}_{l}}\left(\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\right)\ln\frac{n}{|\Delta|}$
$\quad\quad\quad\quad\quad\quad+\frac{1}{|\Delta|}\sum_{z\in\mathcal{Z}_{l}}\left(\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\right)\ln\frac{\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}}{\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}}$
$=\ln\frac{n}{|\Delta|}+\frac{1}{|\Delta|}\sum_{z\in\mathcal{Z}_{l}}\left(\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\right)\ln\frac{\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}}{\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}}$
$\leq\ln\left(\frac{n}{|\Delta|}\right)$
In the third line, we note that $\sum_{z\in\mathcal{Z}_{l}}\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}=|\Delta|$ since each local agent contained in $\Delta$ must have some state/action pair contained in $\mathcal{Z}_{l}$. In the last line, we note that $\sum_{i\in\Delta}\mathbbm{1}\{z_{i}=z\}\leq\sum_{i\in[n]}\mathbbm{1}\{z_{i}=z\}$, For all $z\in\mathcal{Z}_{l}$, and thus the summation of logarithmic terms in the third line is negative. Finally, using this bound in the Bretagnolle-Huber inequality yields the lemma.∎
Corollary E.7.1.
From Lemma E.7, setting $\Delta=[n]$ also recovers $\mathrm{TV}(F_{z_{\Delta}},F_{z_{[n]}})=0$.
Theorem E.8.
With probability atleast $1-\delta$ for $\delta\in(0,1)^{2}$:
$\left|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
Proof.
From combining the total variation distance bound in Lemma E.4 and the Lipschitz continuity bound in Theorem D.3 with $\sum_{t=0}^{T}\gamma^{T}\leq\frac{1}{1-\gamma}$ for $\gamma\in(0,1)$, we have:
$\Pr\left[|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})|\leq\frac{2\epsilon}{1-\gamma}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}\right]\geq 1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$
We then reparameterize $1-2|\mathcal{S}_{l}||\mathcal{A}_{l}|e^{-\frac{8kn\epsilon^{2}}{n-k+1}}$ into $1-\delta$ to get $\epsilon=\sqrt{\frac{n-k+1}{8kn}\ln\left(\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}\right)}$ which gives:
$\Pr\left[\left|\widehat{Q}_{k}^{*}(s_{g},F_{s_{\Delta}},a_{g},F_{a_{\Delta}})-\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},a_{g},F_{a_{[n]}})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}\right]\geq 1-\delta,$
proving the claim.∎
Appendix F Using the Performance Difference Lemma
In general, convergence analysis requires the guarantee that a stationary optimal policy exists. Fortunately, when working with the empirical distribution function, the existence of a stationary optimal policy is guaranteed when the state/action spaces are finite or countably infinite. However, lifting the knowledge of states onto the continuous empirical distribution function space, and designing a policy on the lifted space is still analytically challenging. To circumvent this, Gu et al. (2021) creates lifted $\epsilon$-nets and does kernel regression to obtain convergence guarantees. Here, our analytic approach bears a stark difference, wherein we analyze the sampled structure of the mean-field empirical distribution function, rather than studying the structure of the lifted space. For this, we leverage the classical Performance Difference Lemma, which we restate below for completeness.
Lemma F.1 (Performance Difference Lemma, Kakade & Langford (2002)).
Given policies $\pi_{1},\pi_{2}$, with corresponding value functions $V^{\pi_{1}}$, $V^{\pi_{2}}$:
$V^{\pi_{1}}(s)-V^{\pi_{2}}(s)$ $=\frac{1}{1-\gamma}\mathbb{E}_{\begin{subarray}{c}s^{\prime}\sim d^{\pi_{1}}_{s}\\ a^{\prime}\sim\pi_{1}(\cdot|s^{\prime})\end{subarray}}[A^{\pi_{2}}(s^{\prime},a^{\prime})]$
Here, $A^{\pi_{2}}(s^{\prime},a^{\prime}):=Q^{\pi_{2}}(s^{\prime},a^{\prime})-V^{\pi_{2}}(s^{\prime})$ and $d_{s}^{\pi_{1}}(s^{\prime})=(1-\gamma)\sum_{h=0}^{\infty}\gamma^{h}\Pr_{h}^{\pi_{1}}[s^{\prime},s]$ where $\Pr_{h}^{\pi_{1}}[s^{\prime},s]$ is the probability of $\pi_{1}$ reaching state $s^{\prime}$ at time step $h$ starting from state $s$.
We denote our learned policy ${\pi}_{k,m}^{\mathrm{est}}$ where:
${\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{[n]})=({\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]}),{\pi}_{k,m}^{1,\mathrm{est}}(s_{g},s_{[n]}),\dots,{\pi}_{k,m}^{n,\mathrm{est}}(s_{g},s_{[n]}))\in\mathcal{P}(\mathcal{A}_{g})\times\mathcal{P}(\mathcal{A}_{l})^{n},$
where ${\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{[n]})=\widehat{\pi}_{k,m}^{g,\mathrm{est}}(s_{g},s_{u},F_{s_{\Delta\setminus u}})$ is the global agent’s action and ${\pi}_{k,m}^{i,\mathrm{est}}(s_{g},s_{[n]}):=\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Here, $\Delta_{i}$ is a random variable supported on (${[n]\setminus i\choose k-1}$, $u$ is a random variable uniformly distributed on $[n]$, and $\Delta$ is a random variable uniformly distributed on $[n]\setminus u$. Then, denote the optimal policy $\pi^{*}$ given by
$\pi^{*}(s)=(\pi^{*}_{g}(s_{g},s_{[n]}),\pi_{1}^{*}(s_{g},s_{[n]}),\dots,\pi_{n}^{*}(s_{g},s_{[n]}))\in\mathcal{P}(\mathcal{A}_{g})\times\mathcal{P}(\mathcal{A}_{l})^{n},$
where $\pi^{*}_{g}(s_{g},s_{[n]})$ is the global agent’s action, and $\pi_{i}^{*}(s_{g},s_{[n]}):=\pi^{*}(s_{g},s_{i},F_{s_{\Delta_{i}}})$ is the action of the $i$’th local agent. Next, to compare the difference in the performance of $\pi^{*}(s)$ and ${\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{[n]})$, we define the value function of a policy $\pi$ to be the infinite-horizon $\gamma$-discounted rewards, (denoted $V^{\pi}$) as follows:
Definition F.2.
The value function $V^{\pi}:\mathcal{S}\to\mathbb{R}$ of a given policy $\pi$, for $\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$ is:
$V^{\pi}(s)=\mathbb{E}_{\begin{subarray}{c}a(t)\sim\pi(\cdot|s(t))\end{subarray}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s(t),a(t))|s(0)=s\right].$ (31)
Theorem F.3.
For the optimal policy $\pi^{*}$ and the learned policy ${\pi}_{k,m}^{\mathrm{est}}$, for any state $s_{0}\in\mathcal{S}$, we have:
$V^{\pi^{*}}(s_{0})-V^{{\pi}^{\mathrm{est}}_{k,m}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Proof.
Applying the performance difference lemma to the policies gives us:
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}^{\mathrm{est}}_{k,m}}(s_{0})$ $=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}[V^{\pi^{*}}(s)-Q^{\pi^{*}}(s,a)]$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\bigg{[}\mathbb{E}_{a^{\prime}\sim\pi^{*}(\cdot|s)}Q^{\pi^{*}}(s,a^{\prime})-\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}Q^{\pi^{*}}(s,a)\bigg{]}$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}}\bigg{[}Q^{\pi^{*}}(s,\pi^{*}(\cdot|s))-\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}Q^{\pi^{*}}(s,a)\bigg{]}$
Next, by the law of total expectation,
$\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}\left[Q^{*}(s,a)\right]$
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\sum_{\Delta^{2}\in{[n]\setminus 2\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
Therefore,
$Q^{\pi^{*}}(s,\pi^{*}(\cdot|s))$ $-\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}Q^{*}(s,a)$
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\sum_{\Delta^{2}\in{[n]\setminus 2\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\bigg{(}Q^{*}(s,\pi^{*}(\cdot|s))$
$-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))\bigg{)}$
Therefore, by grouping the equations above, we have:
$V^{\pi^{*}}(s_{0})$ $-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$
$\leq\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim\widetilde{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{i}\in{[n]\setminus i\choose k-1},\\ \forall i\in[n]\end{subarray}}\frac{1}{n{n\choose k}{n-1\choose k-1}^{n}}\bigg{(}\sum_{i\in[n]}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$+\frac{\frac{1}{n}}{{n\choose k}{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}\bigg{)}$
Lemma F.4 shows a uniform bound on
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\Delta^{1}}}),\dots,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{n},F_{s_{\Delta^{n}}}))$
(independent of $\Delta^{i},\forall i\in[n]$), allowing the sums and the counts in the denominator will cancel out.
Observe that $\widehat{\pi}_{k,m}^{*}:\mathcal{S}\to\mathcal{A}$ and $\pi^{*}:\mathcal{S}\to\mathcal{A}$ are deterministic functions. Therefore, denote $a=\pi^{*}(s)$. Then, from Lemma F.6, we have that
$\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\mathbb{E}_{a\sim{\pi}_{k,m}^{\mathrm{est}}(\cdot|s)}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{1}{n{n\choose k}{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}Q^{*}(s,\pi^{*}(s))$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{n}^{*}(s_{g},F_{s_{[n]}},\widehat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{g},\widehat{\pi}_{n}^{*}(s_{g},F_{s_{[n]}})_{1:n})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Similarly, from Corollary F.6.1, we have that
$\frac{1}{1-\gamma}\mathbb{E}_{s\sim d_{s_{0}}^{{\pi}_{k,m}^{\mathrm{est}}}}\sum_{\Delta\in{[n]\choose k}}\sum_{{\Delta^{1}\in{[n]\setminus 1\choose k-1}}}\dots\sum_{{\Delta^{n}\in{[n]\setminus n\choose k-1}}}\frac{\frac{1}{n{n\choose k}}}{{n-1\choose k-1}^{n}}\bigg{|}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Hence, combining the above inequalities, we get:
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{2\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
which yields the claim. We defer parameter optimization to Lemma F.7.∎
Lemma F.4 (Uniform Bound on $Q^{*}$ with different actions).
For all $s\in S,\Delta\in{[n]\choose k}$ and $\Delta^{i}\in{[n]\setminus i\choose k-1}$,
$Q^{*}(s,\pi^{*}(\cdot|s))-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$+\left|\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\right|$
Proof.
Observe that
$Q^{*}(s,\pi^{*}(\cdot|s))$ $-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}})\}_{i\in[n]})$
$\leq\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}}),\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\Delta^{i}})$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\quad\quad-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{k}})$
$\leq\left|Q^{*}(s,\pi^{*}(s))-\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\pi^{*}(s)_{i},\{\pi^{*}(s)_{j}\}_{j\in\Delta^{i}})\right|$
$\quad\quad+\bigg{|}\frac{1}{n}\sum_{i\in[n]}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|}$
$\leq\frac{1}{n}\sum_{i\in[n]}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad+\frac{1}{n}\sum_{i\in[n]}\bigg{|}\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j,\Delta^{j}})\}_{j\in\{i,\Delta^{i}\}})$
$\quad\quad\quad\quad-Q^{*}(s,\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})_{g},\{\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{s_{\Delta^{j}}})\}_{j\in[n]}\bigg{|},$
which proves the claim.∎
Lemma F.5.
Fix $s\in\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$. For each $j\in[n]$, suppose we are given $T$-length sequences of random variables $\{\Delta^{j}_{1},\dots,\Delta^{j}_{T}\}_{j\in[n]}$, distributed uniformly over the support (${[n]\setminus j\choose k-1}$. Further, suppose we are given a fixed sequence $\delta_{1},\dots,\delta_{T}$, where each $\delta_{t}\in(0,1]$ for $t\in[T]$. Then, for each action $a=(a_{g},a_{[n]})=\pi^{*}(s)$, for $t\in[T]$ and $j\in[n]$, define deviation events $B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}$ such that:
$B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}\!:=\!\left\{\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{j},a_{g})\!-\!\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{j},a_{g})\right|\!>\!\frac{\sqrt{\frac{n-k+1}{8kn}\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}}{1-\gamma}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}+\epsilon_{k,m}\right\}$ (32)
For $i\in[T]$, we define bad-events $B_{t}$ (which unifies the deviation events) such that
$B_{t}=\bigcup_{j\in[n]}\bigcup_{\begin{subarray}{c}a_{g}\in\mathcal{A}_{g}\end{subarray}}\bigcup_{a_{j,\Delta_{t}^{j}}\in\mathcal{A}_{l}^{k}}B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}$
Next, denote $B=\cup_{i=1}^{T}B_{i}$. Then, the probability that no bad event $B_{t}$ occurs is:
$\Pr\left[\bar{B}\right]\coloneq 1-\Pr[B]\geq 1-|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\sum_{i=1}^{T}\delta_{i}$
Proof.
$|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})$ $-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})|$
$=\bigg{|}{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})$
$\quad\quad+\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\bigg{|}$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\quad\quad+\left|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$
$\leq\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|+\epsilon_{k,m}$
The first inequality above follows from the triangle inequality, and the second inequality uses
$\left|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|$ $\leq\left\|\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right\|_{\infty}$
$\leq\epsilon_{k,m},$
where the $\epsilon_{k,m}$ follows from Lemma 3.2. From Theorem E.8, we have that with probability at least $1-\delta_{t}$,
$\left|{Q}^{*}(s_{g},s_{j},F_{z_{[n]\setminus j}},a_{g},a_{j})-\widehat{Q}_{k}^{*}(s_{g},s_{j},F_{z_{\Delta_{t}^{j}}},a_{g},a_{j})\right|\leq\frac{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta_{t}}}{1-\gamma}\sqrt{\frac{n-k+1}{8kn}}\cdot\|r_{l}(\cdot,\cdot)\|_{\infty}$
So, event $B_{t}^{a_{g},a_{j,\Delta_{t}^{j}}}$ occurs with probability atmost $\delta_{t}$.
Here, observe that if we union bound across all the events parameterized by the empirical distributions of $a_{g},a_{j,\Delta_{t}^{j}}\in\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ given by $F_{a_{\Delta_{t}^{j}}}$, this also forms a covering of the choice of variables $F_{s_{\Delta_{t}^{j}}}$ by agents $j\in[m]$, and therefore across all choices of $\Delta_{t}^{1},\dots,\Delta_{t}^{n}$ (subject to the permutation invariance of local agents) for a fixed $t$. Thus, from the union bound, we get:
$\Pr[B_{t}]\leq\sum_{a_{g}\in\mathcal{A}_{g}}\sum_{F_{a_{\Delta_{t}}}\in\mu_{k}(\mathcal{A}_{l})}\Pr[B_{t}^{a_{g},a_{1,\Delta_{t}^{1}}}]\leq|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\Pr[B_{t}^{a_{g},a_{1,\Delta_{t}^{1}}}]$
Applying the union bound again proves the lemma:
$\Pr[\bar{B}]$ $\geq 1-\sum_{t=1}^{T}\Pr[B_{t}]\geq 1-|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\sum_{t=1}^{T}\delta_{t}$
Lemma F.6.
For any arbitrary distribution $\mathcal{D}$ of states $\mathcal{S}:=\mathcal{S}_{g}\times\mathcal{S}_{l}^{n}$, for any $\Delta^{i}\in{[n]\setminus i\choose k-1}$ for $i\in[n]$ and for $\delta\in(0,1]$ we have:
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Proof.
By the linearity of expectations, observe that:
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$=\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
Then, define the indicator function $\mathcal{I}:[n]\times\mathcal{S}\times\mathbb{N}\times(0,1]\to\{0,1\}$ by:
$\mathcal{I}(i,s,k,\delta):=$
$\quad\mathbbm{1}\left\{\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\leq\frac{\|r_{l}(\cdot,\cdot)\|_{\infty}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}\right\}$
We then study the expected difference between $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))$ and $\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})$.
Observe that:
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad\quad\quad\quad\quad=\mathbb{E}_{s\sim\mathcal{D}}\left[\mathcal{I}(i,s,k,\delta)\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s\sim\mathcal{D}}\left[(1-\mathcal{I}(i,s,k,\delta))\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|\right]$
Here, we have used the general property for a random variable $X$ and constant $c$ that $\mathbb{E}[X]=\mathbb{E}[X\mathbbm{1}\{X\leq c\}]+\mathbb{E}[(1-\mathbbm{1}\{X\leq c\})X]$. Then,
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})\right|$
$\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}\left(1-\mathbb{E}_{s^{\prime}\sim\mathcal{D}}\mathcal{I}(s^{\prime},k,\delta)\right))$
$\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$\quad=\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
For the first term in the first inequality, we use $\mathbb{E}[X\mathbbm{1}\{X\leq c\}]\leq c$. For the second term, we trivially bound $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k}^{*}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})$ by the maximum value $Q^{*}$ can take, which is $\frac{\widetilde{r}}{1-\gamma}$ by Lemma B.5. In the second inequality, we use the fact that the expectation of an indicator function is the conditional probability of the underlying event. The second inequality follows from Lemma F.5.
Since this is a uniform bound that is independent of $\Delta^{j}$ for $j\in[n]$ and $i\in[n]$, we have:
$\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad\quad\quad\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
This proves the claim.∎
Corollary F.6.1.
Changing the notation of the policy to ${\pi}_{k,m}^{\mathrm{est}}$ in the proofs of Lemmas F.6 and F.5,such that $a\sim\widetilde{\pi}_{k,m}^{\mathrm{est}}$ then yields that:
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\sum_{\Delta^{2}\in{[n]\setminus 2\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,a)-\widehat{Q}^{*}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},a_{g},a_{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Lemma F.7 (Optimizing Parameters).
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)$
Proof.
Setting $\delta=\frac{(1-\gamma)^{2}}{2\widetilde{r}\varepsilon|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}$ in the bound for the optimality gap in Theorem F.3 gives:
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$=\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{4\widetilde{r}\varepsilon|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}{(1-\gamma)^{2}}}+\frac{1}{\varepsilon}+2\epsilon_{k,m}$
Setting $\varepsilon=10\sqrt{k}$ for any $c>0$ recovers a decaying optimality gap of the order
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\widetilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
Finally, setting $\epsilon_{k,m}\leq\widetilde{O}(\frac{1}{\sqrt{k}})$ from Lemma F.9 yields that
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right),$
which proves the claim.∎
F.1 Proof of Lemma 3.2: Bounding the Bellman Error
Theorem F.8 (Theorem 2 of Li et al. (2022)).
If $m\in\mathbb{N}$ is the number of samples in the Bellman update, there exists a universal constant $20<c_{0}<$ 2 and a Bellman noise $0<\epsilon_{k,m}\leq\frac{1}{1-\gamma}$ such that
$\|\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{*}\|_{\infty}=\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\leq\epsilon_{k,m},$
where $\epsilon_{k,m}$ satisfies
$m=\frac{c_{0}\cdot\widetilde{r}\cdot t_{\text{cover}}}{(1-\gamma)^{5}\epsilon_{k,m}^{2}}\log^{2}(|\mathcal{S}||\mathcal{A}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right),$ (33)
where $t_{\text{cover}}$ stands for the cover time, and is the time taken for the trajectory to visit all state-action pairs at least once. Formally,
$t_{\text{cover}}\coloneq\min\left\{t\bigg{|}\min_{(s_{0},a_{0})\in\mathcal{S}\times\mathcal{A}}\mathbb{P}(\mathcal{B}_{t}|s_{0},a_{0})\geq\frac{1}{2}\right\},$ (34)
where $\mathcal{B}_{t}$ denotes the event that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ have been visited at least once between time 00 and time $t$, and $\mathbb{P}(\mathcal{B}_{t}|s_{0},a_{0})$ denotes the probability of $\mathcal{B}_{t}$ conditioned on the initial state $(s_{0},a_{0})$.
Lemma F.9.
The SUB-SAMPLE-MFQ: Learning algorithm runs in time $\widetilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{2.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\widetilde{r})$, while attaining a Bellman noise $\epsilon_{k,m}$ of $\widetilde{O}(1/\sqrt{k})$.
Proof.
Rearranging Eq. 33, one has
$\epsilon_{k,m}\leq\frac{k\sqrt{2\widetilde{r}t_{\text{cover}}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (35)
Then, using the naïve bound $t_{\text{cover}}$ (since we are doing offline learning), we have
$t_{\text{cover}}\leq|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}.$
Substituting this in Eq. 35 gives
$\epsilon_{k,m}\leq\frac{k\sqrt{2\widetilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (36)
Therefore, setting
$m=\frac{2\widetilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{2.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}}{(1-\gamma)^{5}}\log(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (37)
attains a Bellman error of $\epsilon_{k,m}\leq\widetilde{O}(1/\sqrt{k})$. Finally, the runtime of our learning algorithm is
$O(m|\mathcal{S}_{g}||\mathcal{S}_{l}||\mathcal{A}_{g}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|})=\widetilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{2.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\widetilde{r}),$
which is still polynomial in $k$, proving the claim.∎
Appendix G Generalize to Stochastic Rewards
Suppose we are given two families of distributions, $\{\mathcal{G}_{s_{g},a_{g}}\}_{s_{g},a_{g}\in\mathcal{S}_{g}\times\mathcal{A}_{g}}$ and $\{\mathcal{L}_{s_{i},s_{g},a_{i}}\}_{s_{i},s_{g},a_{i}\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}$. Let $R(s,a)$ denote a stochastic reward of the form
$R(s,a)=r_{g}(s_{g},a_{g})+\sum_{i\in[n]}r_{l}(s_{i},s_{g},a_{i}),$ (38)
where the rewards of the global agent $r_{g}$ emerge from a distribution $r_{g}(s_{g},a_{g})\sim\mathcal{G}_{s_{g},a_{g}}$, and the rewards of the local agents $r_{l}$ emerge from a distribution $r_{l}(s_{i},s_{g},a_{i})\sim\mathcal{L}_{s_{i},s_{g},a_{i}}$. For $\Delta\subseteq[n]$, let $R_{\Delta}(s,a)$ be defined as:
$R_{\Delta}(s,a)=r_{g}(s_{g},a_{g})+\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})$ (39)
We make some standard assumptions of boundedness on $\mathcal{G}_{s_{g},a_{g}}$ and $\mathcal{L}_{s_{i},s_{g},a_{i}}$.
Assumption G.1.
Define
$\bar{\mathcal{G}}=\bigcup_{(s_{g},a_{g})\in\mathcal{S}_{g}\times\mathcal{A}_{g}}\mathrm{supp}\left(\mathcal{G}_{s_{g},a_{g}}\right),\quad\quad\bar{\mathcal{L}}=\bigcup_{(s_{i},s_{g},a_{i})\in\mathcal{S}_{l}\times\mathcal{S}_{g}\times\mathcal{A}_{l}}\mathrm{supp}\left(\mathcal{L}_{s_{i},s_{g},a_{i}}\right),$ (40)
[…]
, and
=inf(¯)
infimum¯
[…]
are all known in advance.
Definition G.2.
Let the randomized empirical adapted Bellman operator be $\widehat{\mathcal{T}}^{\text{random}}_{k,m}$ such that:
$\widehat{\mathcal{T}}^{\text{random}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\Delta}},a_{g},a_{1})=R_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{\ell\in[m]}\max_{a^{\prime}\in\mathcal{A}}\widehat{Q}_{k,m}^{t}(s_{g}^{\ell},s_{j}^{\ell},F_{z_{\Delta\setminus j}^{\ell}},a_{j}^{\ell},a_{g}^{\ell}),$ (41)
SUBSAMPLE-MFQ: Learning with Stochastic Rewards.
The proposed algorithm averages $\Xi$ samples of the adapted randomized empirical adapted Bellman operator, $\mathcal{T}_{k,m}^{\text{random}}$ and updates the $\widehat{Q}_{k,m}$ function using with the average. One can show that $\mathcal{T}_{k,m}^{\text{random}}$ is a contraction operator with module $\gamma$. By Banach’s fixed point theorem, $\mathcal{T}_{k,m}^{\text{random}}$ admits a unique fixed point $\widehat{Q}_{k,m}^{\text{random}}$.
Algorithm 5 SUBSAMPLE-MFQ: Learning with Stochastic Rewards
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
1:  Set $\widetilde{\Delta}=\{2,\dots,k\}$.
2:  Set $\mu_{k-1}(Z_{l})=\{0,\frac{1}{k-1},\frac{2}{k-1},\dots,1\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$.
3:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})=0$, for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}$.
4:  for $t=1$ to $T$ do
5:     for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}\times\mathcal{A}_{g}\times\mathcal{A}_{l}$ do
6:        $\rho=0$
7:        for $\xi\in\{1,\dots,\Xi\}$ do
8:           $\rho=\rho+\widehat{\mathcal{T}}^{\text{random}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})$
9:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{g},a_{1})=\rho/\Xi$
10:  For all $(s_{g},s_{i},F_{s_{\Delta\setminus i}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}{(\mathcal{S}_{l})}$, let
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta\setminus i}})=\arg\max_{(a_{g},a_{i},F_{a_{\Delta\setminus i}})\in\mathcal{A}_{g}\times\mathcal{A}_{l}\times\mu_{k-1}{(\mathcal{Z}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{i},F_{z_{\Delta\setminus i}},a_{g},a_{i}).$
Theorem G.3 (Hoeffding’s Theorem (Tsybakov, 2008)).
Let $X_{1},\dots,X_{n}$ be independent random variables such that $a_{i}\leq X_{i}\leq b_{i}$ almost surely. Then, let $S_{n}=X_{1}+\dots+X_{n}$. Then, for all $\epsilon>0$,
$\mathbb{P}[|S_{n}-\mathbb{E}[S_{n}]|\geq\epsilon]\leq 2\exp\left(-\frac{2\epsilon^{2}}{\sum_{i=1}^{n}(b_{i}-a_{i})^{2}}\right)$ (42)
Lemma G.4.
When ${\pi}^{\mathrm{est}}_{k,m}$ is derived from the randomized empirical value iteration operation and applying our online subsampling execution in Algorithm 3, we get
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}}.$ (43)
Proof.
[…]
2
[…]
2
[…]
2
[…]
2
[…]
Rearranging this, we get:
[…]
2
2Ξ2
1δ
[…]
2
2superscriptΞ2
[…]
k1/4ln(200k)
[…]
gives:
$\Pr\left[\left|\frac{\rho}{\Xi}-\mathbb{E}[R_{\Delta}(s,a)]\right|\leq\frac{1}{\sqrt{200k}}\right]\geq 1-\frac{1}{100\sqrt{k}}$
Then, applying the triangle inequality to $\epsilon_{k,m}$ allows us to derive a probabilistic bound on the optimality gap between $\widehat{Q}_{k,m}^{\mathrm{est}}$ and $Q^{*}$, where the gap is increased by $\frac{1}{\sqrt{200k}}$, and where the randomness is over the stochasticity of the rewards. Then, the optimality gap between $V^{\pi*}$ and $V^{{\pi}_{k,m}^{\mathrm{est}}}$, for when the policy $\widehat{\pi}^{\mathrm{est}}_{k,m}$ is learned using Algorithm 5 in the presence of stochastic rewards obeying G.1 follows
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)\right]\geq 1-\frac{1}{100\sqrt{k}},$ (46)
which proves the lemma.∎
Remark G.5.
First, note that that through the naïve averaging argument, the optimality gap above still decays to 00 with probability decaying to $1$, as $k$ increases.
Remark G.6.
[…]
and using methods from order statistics to bound the errors of the estimates and use the deviations between estimates to determine an optimal online stopping time (Kleinberg, 2005) as is done in the online secretary problem. This, more sophisticated, argument would become an essential step in converting this to a truly online learning, with a stochastic approximation scheme. Furthermore, one could incorporate variance-based analysis in the algorithm, where we use information derived from higher-order moments to do a weighted update of the Bellman update (Jin et al., 2024), thereby taking advantage of the skewedness of the distribution, and allowing us the freedom to assign an optimism/pessimism score to our estimate of the reward.
Appendix H Extension to Continuous State/Action Spaces
Preliminaries.
For a measurable space $(\mathcal{S},\mathcal{B})$, where $\mathcal{B}$ is a $\sigma$-algebra on $\mathcal{S}$, let $\mathbb{R}^{\mathcal{S}}$ denote the set of all real-valued $\mathcal{B}$-measurable functions on $\mathcal{S}$. Let $(X,d)$ be a metric space, where $X$ is a set and $d$ is a metric on $X$. A set $S\subseteq X$ is dense in $X$ if every element of $X$ is either in $S$ or a limit point of $S$. A set is nowhere dense in $X$ if the interior of its closure in $X$ is empty. A set $S\subseteq X$ is of Baire first category if $S$ is a union of countably many nowhere dense sets. A set $S\subseteq X$ is of Baire second category if $X\setminus S$ is of first category.
Theorem H.1 (Baire Category Theorem (Jin et al., 2020)).
Let $(X,d)$ be a complete metric space. Then any countable intersection of dense open subsets of $X$ is dense.
Definition H.2.
$\mathrm{MDP}(S,A,\mathbb{P},r)$ is a linear MDP with feature map $\phi:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$ if there exist $d$ unknown (signed) measures $\mu=(\mu^{1},\dots,\mu^{d})$ over $\mathcal{S}$ and a vector $\theta\in\mathbb{R}^{d}$ such that for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have
$\mathbb{P}(\cdot|s,a)=\langle\phi(s,a),\mu(\cdot)\rangle,\quad\quad r(s,a)=\langle\phi(s,a),\theta\rangle$ (47)
Without loss of generality, assume $\|\phi(s,a)\|\leq 1,\forall(s,a)\in S\times A$ and $\max\{\|\mu(S)\|,\|\theta\|\}\leq\sqrt{d}$.
We motivate our analysis by reviewing representation learning in RL via spectral decompositions. For instance, if $P(s^{\prime}|s,a)$ admits a linear decomposition in terms of some spectral features $\phi(s,a)$ and $\mu(s^{\prime})$, then the $Q(s,a)$-function can be linearly represented in terms of the spectral features $\phi(s,a)$.
Assumption H.3.
Suppose $\mathcal{S}_{g}\subset\mathbb{R}^{\sigma_{g}},\mathcal{S}_{l}\subset\mathbb{R}^{\sigma_{l}},\mathcal{A}_{g}\subset\mathbb{R}^{\alpha_{g}},\mathcal{A}_{l}\subset\mathbb{R}^{\alpha_{l}}$ are bounded compact sets. From the Baire category theorem, the underlying field $\mathbb{R}$ can be replaced to any set of Baire first category, which satisfies the property that there exists a dense open subset. In particular, replace 2.3 with a boundedness assumption.
These multi-agent settings where each agent handles a continuous state/action space find many applications in optimization, control, and synchronization.
Example H.4 (Federated Learning).
Consider a peer-to-peer learning setting (Chaudhari et al., 2024) with $n$ agents, where each agent possesses a common neural network architecture $f(\mathbf{x};\mathbf{\theta}):\mathbb{R}^{N}\to\mathbb{R}^{M}$, parameterized by $\theta\in\mathbb{R}^{P}$. Here, each local agent has a common local objective function $L_{q}(\mathbf{\theta}):\mathbb{R}^{P}\to\mathbb{R}$ defined using a local dataset $\mathcal{D}_{q}\stackrel{{\scriptstyle\Delta}}{{=}}\{x_{q_{i}},y_{q,i}\}_{i=1}^{D}$. For instance, $L_{q}(\mathbf{\theta})$ can be the local empirical risk function that evaluates the performance of $f(\mathbf{x};\mathbf{\theta})$ on a task involving the local dataset $\mathcal{D}_{q}$, where the agents cooperate to learn the optimal $\mathbf{\theta}^{*}$ that minimizes the global objective, which is a weighted average of all local objectives. Formally, the agents collaboratively aim to find $\mathbf{\theta}^{*}$ satisfying:
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{P}}L(\theta)$
$L(\theta)=\frac{1}{Q}\sum_{q=1}^{Q}L_{q}(\theta)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{D}\sum_{i=1}^{D}\ell(f(x_{q,i};\theta),y_{q,i})$
Here, the global agent may act to balance the loss among agents, by ensuring that the variance on the losses across the agents is small. If the variance on the losses of the agents is large, it could assign a large cost to the system, and signal that the policies of the local agents must favor the mean state, promoting convergence.
Example H.5 (Quadcopter Swarm Drone (Preiss et al., 2017)).
Consider a system of drones with a global controller, where each drone has to chase the controller, and the controller is designed to follow a bounded trajectory. Here, the state of each local agent $i\in[n]$ is its position and velocity in the bounded region, and the state of the global agent $g$ is a signal on its position and direction. The action of each local agent is a velocity vector $a_{l}\in\mathcal{A}_{l}\subset\mathbb{R}^{3}$ which is a bounded subset of $\mathbb{R}^{3}$, and the action of the global agent is a vector $a_{g}\in\mathcal{A}_{g}\subset\mathbb{R}^{2}$ which is a bounded subset of $\mathbb{R}^{2}$.
Lemma H.6 (Proposition 2.3 in Jin et al. (2020)).
For any linear MDP, for any policy $\pi$, there exist weights $\{w^{\pi}\}$ such that for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, we have $Q^{\pi}(s,a)=\langle\phi(s,a),\mathbf{w}^{\pi}\rangle$.
Lemma H.7 (Proposition A.1 in Jin et al. (2020)).
For any linear MDP, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, and for any measurable subset $\mathcal{B}\subseteq\mathcal{S}$, we have that
$\phi(s,a)^{\top}\mu(\mathcal{S})=1,\quad\quad\phi(s,a)^{\top}\mu(\mathcal{B})\geq 0.$
Property H.8.
Suppose that there exist spectral representation $\mu_{g}(s_{g}^{\prime})\in\mathbb{R}^{d}$ and $\mu_{l}(s_{i}^{\prime})\in\mathbb{R}^{d}$ such that the probability transitions $P_{g}(s_{g}^{\prime},|s_{g},a_{g})$ and $P_{l}(s_{i}^{\prime}|s_{i},s_{g},a_{i})$ can be linearly decomposed by
$P_{g}(s_{g}^{\prime}|s_{g},a_{g})=\phi_{g}(s_{g},a_{g})^{\top}\mu_{g}(s_{g}^{\prime})$ (48)
$P_{l}(s_{i}^{\prime}|s_{i},s_{g},a_{i})=\phi_{l}(s_{i},s_{g},a_{i})^{\top}\mu_{l}(s_{i}^{\prime})$ (49)
for some features $\phi_{g}(s_{g},a_{g})\in\mathbb{R}^{d}$ and $\phi_{l}(s_{i},s_{g},a_{i})\in\mathbb{R}^{d}$. Then, the dynamics are amenable to the following spectral factorization, as in Ren et al. (2024).
Lemma H.9.
$\widehat{Q}_{k}$ admits a linear representation.
Proof.
Given the factorization of the dynamics of the multi-agent system, we have:
$\mathbb{P}(s^{\prime}|s,a)$ $=P_{g}(s_{g}^{\prime}|s_{g},a_{g})\cdot\prod_{i=1}^{n}P_{l}(s_{i}^{\prime}|s_{i},s_{g},a_{i})$
$=\langle\phi_{g}(s_{g},a_{g}),\mu_{g}(s_{g}^{\prime})\rangle\cdot\prod_{i=1}^{n}\langle\phi_{l}(s_{i},s_{g},a_{i}),\mu_{l}(s_{i}^{\prime})\rangle$
$=\phi_{g}(s_{g},a_{g}),\mu_{g}(s_{g}^{\prime})\rangle\cdot\left\langle\otimes_{i=1}^{n}\phi_{l}(s_{i},s_{g},a_{i}),\otimes_{i=1}^{n}\mu_{l}(s_{i}^{\prime})\right\rangle\coloneq\langle\bar{\phi}_{n}(s,a),\bar{\mu}_{n}(s^{\prime})\rangle$
Similarly, for any $\Delta\subseteq[n]$ where $|\Delta|=k$, the subsystem consisting of $k$ local agents $\Delta$ has a subsystem dynamics given by
$\mathbb{P}(s_{\Delta}^{\prime},s_{g}|s_{\Delta},s_{g},a_{g},a_{\Delta})=\langle\bar{\phi}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta}),\bar{\mu}_{k}(s_{\Delta}^{\prime})\rangle.$
Therefore, $\widehat{Q}_{k}$ admits the linear representation:
$\widehat{Q}_{k}^{\pi_{k}}(s_{g},F_{z_{\Delta}},a_{g})$ $=r_{g}(s_{g},a_{g})+\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})+\gamma\mathbb{E}_{s_{g}^{\prime},s_{\Delta^{\prime}}}\max_{a_{g}^{\prime},a_{\Delta}^{\prime}}\widehat{Q}_{k}^{\pi_{k}}(s_{g}^{\prime},F_{z_{\Delta}^{\prime}},a_{g}^{\prime})$
$=r_{g}(s_{g},a_{g})+\frac{1}{k}\sum_{i\in\Delta}r_{l}(s_{i},s_{g},a_{i})+(\mathbb{P}V^{\pi_{k}}(s_{g},F_{s_{\Delta}},a_{g}))$
$=\begin{bmatrix}r_{\Delta}(s,a)\\ \bar{\phi}_{k}(s_{g},s_{\Delta},a_{g},a_{\Delta})\end{bmatrix}\begin{bmatrix}1&\gamma\int_{s_{\Delta}^{\prime}}\mu_{k}(s_{\Delta}^{\prime})V^{\pi_{k}}(s_{\Delta}^{\prime})\mathrm{d}s_{\Delta}^{\prime}\end{bmatrix},$
proving the claim. ∎
Therefore, $\bar{\phi}_{k}$ serves as a good representation for the $\widehat{Q}_{k}$-function making the problem amenable to the classic linear functional approximation algorithms, consisting of feature generation and policy gradients.
In feature generation, we generate the appropriate features $\phi$, comprising of the local reward functions and the spectral features coming from the factorization of the dynamics. In applying the policy gradient, we perform a gradient step to update the local policy weights $\theta_{i}$ and update the new policy. For this, we update the weight $w$ via the TD(0) target semi-gradient method (with step size $\alpha$), to get
$w_{t+1}=w_{t}+\alpha(r_{\Delta}(s^{t},a^{t})+\gamma\bar{\phi}_{k}(s^{t+1}_{\Delta})^{\top}w_{t}-\bar{\phi}_{k}(s_{\Delta}^{t})^{\top}w_{t})\bar{\phi}_{k}(s_{\Delta}^{t})\$ (50)
Definition H.10.
Let the complete feature map be denoted by $\Phi\in\mathbb{R}^{|\mathcal{S}|\times d}$, where
$\Phi=\begin{bmatrix}\phi(1)\\ \phi(2)\\ \vdots\\ \phi(S)\end{bmatrix}\in\mathbb{R}^{S\times d}$ (51)
Similarly, let the subsampled feature map be denoted by $\widehat{\Phi}_{k}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}$, where
$\widehat{\Phi}_{k}=\begin{bmatrix}\bar{\phi}_{k}(1)\\ \bar{\phi}_{k}(2)\\ \vdots\\ \bar{\phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|)\end{bmatrix}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}\$ (52)
Here, the system’s stage rewards are given by :
$r=\begin{bmatrix}r(1)&\dots&r(S)\end{bmatrix}\in\mathbb{R}^{S},$ (53)
$r_{k}=[r_{k}(1),\dots,r_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|]\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|},$ (54)
where $d$ is a low-dimensional embedding we wish to learn.
Agnostically, the goal is to approximate the value function through the approximation
$V_{w}\approx\widehat{\Phi}_{k}w=\begin{bmatrix}\widehat{\Phi}_{k}(1)\\ \vdots\\ \widehat{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})})\end{bmatrix}w\in\mathrm{span}(\widehat{\Phi}_{k})$
This manner of updating the weights can be considered via the projected Bellman equation $\widehat{\Phi}_{k}w=\Pi_{\mu}\mathcal{T}(\widehat{\Phi}_{k}w)$, where $\Pi_{\mu}(v)=\arg\min_{z\in\widehat{\Phi}_{k}w}\|z-v\|_{\mu}^{2}$. Notably, the fixed point of the the projected Bellman equation satisfies $w=(\widehat{\Phi}_{k}^{\top}D_{\mu}\widehat{\Phi}_{k})^{-1}\widehat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\widehat{\Phi}_{k}w)$, where $D_{\mu}$ is a diagonal matrix comprising of $\mu$’s. In other words, $D_{\mu}=\mathrm{diag}(\mu_{1},\dots,\mu_{n})$. Then,
$(\widehat{\Phi}_{k}^{\top}D_{\mu}\widehat{\Phi}_{k})w=\widehat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\widehat{\Phi}_{k}w)$
In turn, this implies
$\widehat{\Phi}_{k}^{\top}D_{\mu}(I-\gamma P^{\pi})\Phi w=\widehat{\Phi}_{k}^{\top}D_{\mu}r.$
Therefore, the problem is amenable to Algorithm 1 in Ren et al. (2024). To bound the error of using linear function approximation to learn $\widehat{Q}_{k}$, we use a result from Ren et al. (2024).
Lemma H.11 (Policy Evaluation Error, Theorem 6 of Ren et al. (2024)).
Suppose the sample size $M\geq\log\left(\frac{4n}{\delta}\right)$, where $n$ is the number of agents and $\delta\in(0,1)$ is an error parameter. Then, with probability at least $1-2\delta$, the ground truth $\widehat{Q}^{\pi}_{k}(s,a)$ function and the approximated $\widehat{Q}_{k}$-function $\widehat{Q}^{\mathrm{LFA}}_{k}(s,a)$ satisfies, for any $(s,a)\in\mathcal{S}\times\mathcal{A},$
[…]
where $\epsilon_{P}$ is the error in approximating the spectral features $\phi_{g},\phi_{l}$.
Corollary H.11.1.
Therefore, when ${\pi}_{k,m}$ is derived from the spectral features learned in $\widehat{Q}^{\mathrm{LFA}}_{k}$, applying the triangle inequality on the Bellman noise and setting $\delta=\epsilon_{P}=1/\sqrt{k}$ yields [^1] Following Ren et al. (2024), the result easily generalizes to any positive-definite transition Kernel noise (e.g. Laplacian, Cauchy, Matérn, etc.
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\log\left(2k^{2}\right)\frac{\|\bar{\phi}_{k}\|^{5}}{\sqrt{M}}+\frac{2}{\sqrt{k}}\cdot\frac{\gamma\widetilde{r}}{1-\gamma}\|\bar{\phi}_{k}\|\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\left(1-\frac{2}{\sqrt{k}}\right)$ (55)
Using $\|\widetilde{\phi}_{k}\|\leq 1$ gives
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\frac{\log\left(2k^{2}\right)}{\sqrt{M}}+\frac{2\gamma\widetilde{r}}{\sqrt{k}(1-\gamma)}\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\left(1-\frac{2}{\sqrt{k}}\right)$
Remark H.12.
Hence, in the continuous state/action space setting, as $k\to n$ and $M\to\infty$, $V^{\widetilde{\pi}_{k}}(s_{0})\to V^{\widetilde{\pi}_{n}}(s_{0})=V^{\pi^{*}}(s_{0})$. Intuitively, as $k\to n$, the optimality gap diminishes following the arguments in Corollary F.6.1. As $M\to\infty$, the number of samples obtained allows for more effective learning of the spectral features.
Appendix I Towards Deterministic Algorithms: Sharing Randomness
In this section, we propose algorithms RANDOM-SUB-SAMPLE-MFQ and RANDOM-SUB-SAMPLE-MFQ+, which shares randomness between agents through various careful groupings and using shared randomness within each groups. By implementing these algorithms, we derive simulation results, and establish progress towards derandomizing inherently stochastic approximately-optimal policies.
Algorithm 6 (Execution with weakly shared randomness). The local agents are heuristically divided into arbitrary groups of size $k$. For each group $h_{i}$, the $k-1$ local agents sampled are the same at each iteration. The global agent’s action is then the majority global agent action determined by each group of local agents. At each round, this requires $O(\left\lceil n^{2}/k\right\rceil(n-k))$ random numbers.
Algorithm 6 SUB-SAMPLE-MFQ: Execution with weakly shared randomness
0:  A multi-agent system as described in Section 2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2:  Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:  for $t=0$ to $T^{\prime}-1$ do
4:     for $i\in[x]$ do
5:        Let $\Delta_{i}$ be a uniform sample of (${[n]\setminus g_{i}\choose k-1}$, and let $a_{g_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{1:k}$.
6:     Let $a_{g}(t)=\mathrm{majority}\left(\left\{[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{g}\right\}_{i\in[x]}\right)$.
7:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$
8:     Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t)$
9:     $s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
10:     Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
Algorithm 7 (Execution with strongly shared randomness). The local agents are randomly divided in to groups of size $k$. For each group, each agent uses the other agents in the group as the $k-1$ other sampled states. Similarly, the global agent’s action is then the majority global agent action determined by each group of local agents. Here, at each round, this requires $O(\left\lceil n^{2}/k\right\rceil)$ random numbers.
Algorithm 7 SUB-SAMPLE-MFQ: Execution with strongly shared randomness
0:  A multi-agent system as described in Section 2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2:  Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:  for $t=0$ to $T^{\prime}-1$ do
4:     for $i\in[x-1]$ do
5:        Let $a_{h_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{1:k}$.
6:        Let $a^{g}_{h_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{g}$.
7:     Let $\Delta_{\mathrm{residual}}\coloneq{[n]\setminus h_{x}\choose k-(n-\left\lceil n/k\right\rceil k)}$.
8:     Let $a_{h_{x}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{1:n-\left\lceil n/k\right\rceil k}$
9:     Let $a^{g}_{h_{x}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{g}$
10:     Let $a_{g}(t)=\mathrm{majority}\left(\{a_{h_{i}}^{g}(t):i\in[x]\}\right)$.
11:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
12:     Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
The probability of a bad event (the policy causing the $Q$-function to exceed the Lipschitz bound) scales with the $O(n^{k})$ independent random variables, which is polynomially large. Agnostically, some randomness is always needed to avoid periodic dynamics that may occur when the underlying Markov chain is not irreducible. In this case, an adversarial reward function could be designed such that the system does not minimize the optimality gap, thereby penalizing excessive determinism.
This ties into an open problem of interest, which has been recently explored (Larsen et al., 2024). What is the minimum amount of randomness required for SUBSAMPLE-MFQ (or general mean-field or multi-agent RL algorithms) to perform well? Can we derive a theoretical result that demonstrates and balances the tradeoff between the amount of random bits and the performance of the subsampled policy when using the greedy action from the $k$-agent subsystem derived from $\widehat{Q}_{k}^{*}$? We leave this problem as an exciting direction for future research.
Appendix J Numerical Experiments
This section provides simulation results for the examples outlined in Section 2 for Gaussian squeeze and constrained exploration. All numerical experiments were run on a 3-core CPU server equipped with a RAM. We chose a complexity for the parameters of each simulation that was sufficient to emphasize characteristics of the theory, such as the complexity improvement of Algorithms 1 and 2 and the decaying optimality gap.
J.1 Gaussian Squeeze
In this task, $n$ homogeneous agents determine their individual action $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, $a_{i}=\{0,\dots,9\}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. In scenarios of traffic congestion, each agent $i\in[n]$ is a traffic controller trying to send $a_{i}$ vehicles into the main road, where controllers coordinate with each other to avoid congestion, hence avoiding either over-use or under-use, thereby contributing to the entire system. This GS problem serves as an ablation study on the impact of subsampling for MARL.
Specifically, we consider a purely homogeneous subsystem devoid of a global agent to match the mean-field MARL setting in Yang et al. (2018), where the richness of our model setting can still express the Gaussian squeeze task. Instead, we use the global agent to model the global state of the system, given by the number of vehicles in each controller’s lane.
We set $\mathcal{S}_{l}=\{1,\dots,20\}$, $\mathcal{A}_{l}=\{0,\dots,9\}$, and $\mathcal{S}_{g}=\mathcal{S}_{l}$ such that $s_{g}=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}\right\rceil$. The transition functions are given by $s_{i}(t+1)=s_{i}(t)-1\cdot\mathbbm{1}\{s_{i}(t)>s_{g}(t)\}+\mathrm{Ber}(p)$ and $s_{g}(t+1)=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}(t+1)\right\rceil$, where $\mathrm{Ber}(p)$ is a Bernoulli random variable with parameter $p>0$. Finally, the global agent’s reward function is given by $r_{g}(s_{g})=-s_{g}$ and the local agent’s reward function is given by $r_{l}(s_{i},s_{g},a_{i})=4\cdot\mathbbm{1}\{s_{i}>s_{g}\}-2\cdot\mathbbm{1}\{a_{i}>s_{g}\}$. For this task, we ran two simulations: a large-scale simulation with $n=50$ agents, and a small-scale simulation with $n=6$ agents, and provide simulation results below. We observe the exponential improvement in computational complexity in Figure 6, as well as the monotonic improvement in the cumulative reward in Figure 7.
Figure 6: Computation time (in minutes) against sampling parameter $k$.
Figure 7: Cumulative discounted reward as $k$ increases
J.2 Constrained Exploration
Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent wishes to constrain the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect stage rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, by only using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where some shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Therefore, through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while avoiding unnecessary risk or inefficiency.
We set $\mathcal{S}_{g}=\mathcal{S}_{l}=\{0,1,2,3,4,5,6\}\times\{0,1,2,3,4,5,6\}$ and $\mathcal{A}_{l}=\mathcal{A}_{g}=\left\{\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}1\\ 1\end{pmatrix}\right\}$. Then:
$s_{g}^{t+1}(s_{g}^{t},a_{g}^{t})=\Pi_{S_{g}}(s_{g}^{t}+a_{g}^{t})$
$s_{i}^{t+1}(s_{i}^{t},s_{g}^{t},a_{i}^{t})=\Pi_{S_{l}}(s_{i}^{t}+|s_{i}^{t}-s_{g}^{t}|+a_{i}^{t})$
$r_{g}(s_{g}^{t},a_{g}^{t})=10*\mathbbm{1}\{a_{g}^{t}=0\}$
$r_{l}(s_{i}^{t},s_{g}^{t},a_{i}^{t})=\mathbbm{1}\{a_{i}^{t}=0\}+4\cdot\mathbbm{1}\{|s_{i}^{t}-s_{g}^{t}|=1\}-2\cdot|s_{i}^{t}-s_{g}^{t}|$
For this task, we ran a simulation with $n=20$ agents, and provide simulation results in Fig. 8. Again, we observe the monotonic improvements in the cumulative discounted rewards.
Figure 8: Reward optimality gap (log scale) with ${\pi}_{k,m}^{\mathrm{est}}$ running $300$ iterations.